### v1

# Make Me Happier: Evoking Emotions Through Image Diffusion Models

Abstract
Despite the rapid progress in image generation, emotional image editing remains under-explored. The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design. For the first time, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes. To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments. Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations. Furthermore, we conduct human psychophysics experiments and introduce four new evaluation metrics to systematically benchmark all the methods. Experimental results demonstrate that our method surpasses all competitive baselines. Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images. All code, model, and data will be made public.
Figure 1: The generated images evoke a sense of happiness in viewers, contrasting with the negative emotions elicited by the source images. Given a source image that triggers negative emotions (framed in green), our method (Ours) synthesizes a new image that elicits the given positive target emotions (in red), while maintaining the essential elements and structures of the scene. For instance, in the first quadrant where trash on the ground evokes negative emotions, our model performs image editing by eliminating the trash while preserving the outdoor context, including trees and grass, in the same spatial arrangement as the original scene. For comparisons, we include other competitive methods (Sec. 4). Zoom in for better views.
1 Introduction
“I am feeling down. Enhance my room to bring in more excitement."
What we perceive not only entails useful visual information but also evokes profound emotional reactions. Incorporating visual cues that elicit emotions into images holds practical significance in various domains. For instance, in commercials, it amplifies product branding and shapes public opinions, captivating consumer attention. Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia. Despite significant advancements in image generation techniques [6, 27, 35], there has been a notable absence of focus given to emotion conditioning. Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
Emotion-evoked image generation poses a challenge as it requires a deep understanding and identification of the subtle context and semantic elements present in the source images that evoke emotional responses. Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response. For instance, in Figure 1, the presence of a burning forest significantly contributes to feelings of anger. Consequently, in the generated image, the burning forest is substituted with a serene scene featuring a cluster of flowers in the forest, thereby transforming the overall emotional response to awe.
Figure 2: Emotions are influenced by both global and local cues from the source images. We present two image generation examples showcasing the transition from positive to negative emotions in Column 1 (shaded in blue) and two examples from negative to positive emotions in Column 2 (shaded in orange). In each example, the target emotions are in red, and source images are framed in green. Generated results from three methods are presented. Zoom in for better views.
Studies [31, 32] have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement. As also seen in Figure 2, introducing nauseating or fearful elements like fire or ghosts into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments. However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows. In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses. Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli [2, 11], simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
Drawing inspiration from these emotion-associated factors above, we propose our model, EmoEditor, for emotion-evoked image generation. We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties. (1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome. (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions for image editing. (3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer [21, 7, 30], were commonly employed to modify the emotional tone of images. These methods typically require reference emotional images or textual cues as additional inputs. However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses. In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged [12, 3, 24]. However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images. To address this gap, we propose a competitive baseline (Large Model Series), concatenating multiple large language and vision models. This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models. Upon comparison with existing methods through a series of human psychophysics experiments and four new quantitative evaluations, our end-to-end trained EmoEditor achieves outstanding image generation results. It effectively preserves the structural coherence and semantic consistency wtih the source images while eliciting the target emotion of human viewers.
To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions. Unlike existing datasets on visual emotion analysis [18, 17, 34, 20, 31], which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
The main contributions of this paper are highlighted below: (1) We introduce a new and important problem of emotion-evoked image generation. To benchmark all methods, we establish a comprehensive framework for evaluating model performances, incorporating human psychophysics experiments and four newly introduced metrics. (2) We introduce EmoEditor, an emotion-evoked diffusion model, which employs a novel two-branch architecture to integrate global context with local semantic regions of source images that evoke emotional responses. (3) EmoEditor learns emotional cues during end-to-end training, eliminating the need for additional emotion reference images. During inference, EmoEditor generates emotion-evoked images without hand-crafted emotion-conditioned text instructions, while preserving context and structural coherence with the source image. Its performance is superior among all the competitive methods. (4) We curate the first large-scale EmoPair dataset, containing 340,000 image pairs with emotion annotations.
2 Related Work
2.1 Visual Emotion Analysis
Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses through visual information from images or videos. In psychology, Categorical Emotion States (CES) form the basis for emotion classification models, notably exemplified by the 6-category Ekman model [5] and the 8-category Mikels model [18]. Early research focused on manually crafted features such as color and content [17], while recent advancements leverage deep learning models to automatically capture intrinsic global and local emotional cues for more accurate predictions [33, 32]. Despite advancements in visual emotion analysis, there is a limited focus on generating images that evoke specific emotional responses. To address this, we propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation in a unified framework. Unlike Affective Image Filter (AIF) [30] which relies on explicit textual instructions, EmoEditor does not need explicit text guidance on scene modifications for desired emotions.
Exsiting VEA datasets [20, 17, 34] primarily focus on emotion classification, including EmoSet [31], the largest dataset with 120,000 images annotated for emotion labels and six attributes. However, none of them include image pairs with annotated emotion labels and text editing instructions crucial for emotion-evoked image generation. Therefore, we introduce the EmoPair dataset, comprising 340,000 image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
2.2 Image Editing
Color and Style Transfer. Color and style transfer methods like PDF-Transfer[21], Neural Style Transfer[7], and CLIP-Styler[14] use reference images for image manipulation, with outcomes varying based on choices. While colors and styles may influence human emotions, they struggle to manipulate local image regions, limiting their versatility in emotion-evoked generation. In contrast, our EmoEditor balances local and global image manipulations during inference without requiring additional references.
Text-to-Image Editing. Text-to-image editing is a growing area of interest. Previous studies utilize Generative Adversarial Networks (GAN) [8] architectures and leverage CLIP model’s [22] language embedding for image manipulation from detailed text descriptions[19, 1]. Recently, text-to-image diffusion models like Imagic[12], Prompt-to-prompt[10], InstructPix2Pix[3], and DreamBooth[24] have strong capabilities in generating high-quality and diverse images based on text prompts. Despite their impressive results, they often overlook emotional aspects and rely heavily on detailed textual instructions. In contrast, our EmoEditor only needs target emotion and source image inputs. Through end-to-end training, it comprehends emotional cues and creatively edits images for the desired emotions while preserving structure and coherence.
3 Method
We formulate emotion-evoked image generation as a supervised learning task. First, we curate the EmoPair dataset containing image pairs with emotion labels and corresponding text instructions for editing (Sec. 3.1). Next, EmoEditor is trained on this dataset (Sec. 3.2). During inference, it takes source images and target emotions, producing generated images without the need for human-annotated text instructions or emotion reference images.
Figure 3: Pipeline for Curating Our EmoPair Dataset. The dataset comprises two subsets: EmoPair-Annotated Subset (EPAS, left blue box) and EmoPair-Generated Subset (EPGS, right orange box). Each subset includes schematics depicting the creation, selection, and labeling of image pairs in the upper quadrants, with two example pairs in the lower quadrants. Each example pair comprises a source image (framed in green) and a target image. The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided. See more examples in Supp.. See Sec. 3.1 for EmoPair dataset.
3.1 Curating Our EmoPair Dataset
Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair. It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2p[3] annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions. All images from EmoPair are of size 224$\times$224. Similar to existing works in VEA [18, 31], all images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions. Figure 3 shows the pipeline for curating our EmoPair dataset.
EmoPair-Annotated Subset (EPAS). The original Ip2p dataset [3] lacks emotion labels, consisting only of image pairs with text instructions for editing. To address this limitation, we use an emotion predictor $\mathcal{P}$ to classify the source and target images from the Ip2p dataset into eight distinct emotions. $\mathcal{P}$ employs ResNet18 [9] as a backbone and trained on the EmoSet dataset [31]. It achieves a top-1 accuracy of 73% on EmoSet’s test set. We augment the Ip2p dataset with emotion labels, resulting in a total of 331,595 image pairs. Notably, the text instructions with these image pairs serve as human expectations to guide EmoEditor’s training (Sec. 3.2), but are not utilized during inference.
EmoPair-Generated Subset (EPGS). Due to the lack of salient emotional cues in the Ip2p dataset, we further complement EPAS with EPGS subset, which has emotionally rich images from the EmoSet dataset [31]. EmoSet provides diverse and evocative emotional cues that benefit our EmoEditor. However, EmoSet lacks paired images with coherent scene structures and target emotions, along with the absence of text editing instructions from source to target images.
Directly inputting a single-word prompt like “[Target Emotion]" into existing text-to-image editing models proves ineffective due to their limited emotional understanding and reasoning capabilities. Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images. For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape. Thus, to ensure accurate image pair generation, we created 50 general instructions, agnostic of source images, to prompt transitions to desired emotions across 8 categories using GPT-3 [4]. Human annotators then ranked these instructions per emotion category based on efficacy and retained the top ten. Next, the pre-trained Ip2p model manipulated EmoSet source images using these selected instructions to generate target images.
We employ the following selection criteria to control the quality of generated image pairs: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions. (2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) [29] and Learned Perceptual Image Patch Similarity (LPIPS) [36] for filtering the remaining outcomes. SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences. Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1. Ultimately, the EPGS subset retains 6,949 image pairs.
3.2 Our Diffusion Model - EmoEditor
Figure 4: Architecture of Our Proposed EmoEditor. EmoEditor is an image diffusion model, consisting of local (shaded in green) and global (shaded in blue) branches. The pre-trained VAE’s encoder $\mathcal{E}$ and decoder $\mathcal{D}$ remain fixed throughout training and inference. Exclusively employed during inference, the fixed emotion predictor $\mathcal{P}$ predicts the emotions on generated images for the iterative emotion inference. See Sec.3 for details.
Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion. It employs the latent diffusion model (LDM) [23] with three key novelties. EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios. (2) Source images can be complex, conveying multiple emotions simultaneously. Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
Integration of Global and Local Emotion Cues. In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0. Next, we introduce an emotion encoder $\tau_{\theta}$, which encodes $e_{oh}$ to $e$ via a stack of fully connected blocks: $e=\tau_{\theta}(e_{oh})$.
Unlike classical text-to-image models using explicit text prompts, our $\tau_{\theta}$ is trained from scratch, learning neuro-symbolic embeddings for the reverse diffusion process without relying on pre-trained language models like CLIP[22]. Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$. Notably, we intentionally avoid using explicit text editing instructions from image pairs as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions. Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$. Similar to the LDM [23], $\mathcal{E}$ is based on a pre-trained variational autoencoder[13]. The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps and $t\in T$. Specifically, we expressed $z_{t}$ as $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution with mean 0 and variance 1.
In the reverse diffusion process, the denoising network $\epsilon_{\theta}$ predicts noise added to the noisy latent $z_{t}$ using cross-attention with the target emotion condition $e$ and source image condition $z$. The features in $z$ and $z_{t}$ enhance understanding of intrinsic information in the image, leading to emotion-specific content for $I_{s}$. After $T-1$ denoising steps, decoder $\mathcal{D}$ produces an emotion-evoked image.
We minimize the latent diffusion loss, which is the expected error between the predicted noise by $\epsilon_{\theta}$ and the actual noise $\epsilon$ sampled during training: $L_{noise}=E_{t,y,z,\epsilon(z_{t}),\epsilon\sim N(0,1)}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,z,\tau_{\theta}(y))\right\|^{2}$.
Alignment Loss between Model and Human Judgements. A single emotion term lacks the depth to convey the nuanced causes underlying emotions, posing a challenge for diffusion models to grasp the diagnostic information triggering emotions. In contrast to existing text-to-image models, EmoEditor emulates human reasoning by aligning its thought process with text instructions in our EmoPair dataset, which leads the transition from the source emotion to the target emotion. To achieve this, we introduce an alignment loss in a neuro-symbolic space, between emotion embeddings $e$ from $\tau_{\theta}$ and text embeddings $c$ of instructions in each image pair of our EmoPair dataset with pre-trained CLIP [22]. Specifically, the alignment loss is formulated as the inverse of cosine similarity between $e$ and $c$: $L_{emb}=1-\cos(e,c)$. Overall, we conduct joint training of $\tau_{\theta}$ and $\epsilon_{\theta}$ with the following loss: $L_{total}=L_{noise}+\lambda L_{emb}$, where $\lambda=0.5$ represents the weight for balancing different losses.
Importantly, we differentiate our alignment loss from the losses typically employed in classical language models, which focus on predicting specific word tokens. Our approach loosens the constraint of precisely predicting the exact text instructions. This flexibility empowers our model to explore a wider array of image editing solutions, that can augment the source emotions of human viewers to target emotions. Consequently, our model learns to reason in a neuro-symbolic manner akin to humans, and meanwhile, preserves its creativity and capacity for exploring novel editing solutions.
Iterative Emotion Inference. Editing emotions is a complex process often requiring iterative image edits. To address this, we introduce a recurrent emotion critic process during inference. This critic iteratively infers the evoked emotions from generated images while assessing their structural coherence and semantic consistency. Specifically, we employ the emotion predictor $\mathcal{P}$ as the critic.
In the first iteration of image generation, the source image $I_{s}$ serves as both the input and condition image, producing latent variables $z$ and $z_{t}$. These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$. The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.3-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.8. If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images. In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition. Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic. This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 20. In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
Implementation Details. Follow Wang et al.[28], we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p[3] weights. Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$. The frozen emotion predictor $\mathcal{P}$ is only used during inference. We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer. The max time step $T$ is set to 1,000.
4 Experiment
We selected 504 images randomly from the EmoSet dataset [31], ensuring they were distinct from those in our EmoPair dataset. According to the emotion labels of EmoSet, there are 63 images for each source emotion category. Subsequently, we evaluated both competitive baselines and our EmoEditor on these source images and presented the results in the following subsections.
Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions. Considering emotional valence, we divided the test set into two subsets: within-valence emotion transfer and cross-valence emotion transfer. This division serves two purposes. Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween. Secondly, images of the same valence may contain subtle differences challenging for human viewers to discern. For instance, an amusement park scene might evoke a mixture of excitement and amusement.
Baselines. We compare our EmoEditor with five state-of-the-art methods: (1) Color-transfer[21] (CT); (2) Neural-Style-Transfer[7] (NST). Both methods require an additional reference image from the target emotion category as inputs. We randomly selected these reference images from the EmoSet dataset. We also include representative text-to-image editing models (3) CLIP-Styler[14] (Csty) and (4) Ip2p [3]. We provide these models with the target emotion category as a one-word text prompt, such as “excitement". Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS). This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions. See Supp. for details.
Human Psychophysics Experiment. We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)[26]. We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. All participants are properly compensated. In each trial, participants engage in a “two-alternative forced choice" task. They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them. The two image stimuli consists of a result generated by our EmoEditor and a randomly sampled result from the five baselines introduced above. The trial presentation order and the binary choices for each trial are randomized. To assess the quality of data collection, we introduce control trials and use them as filtering criteria. We discard the data from those people who fail the control trials at least 17% of the time. See Supp. for more experimental details.
Evaluation Metrics. We introduce four novel evaluation metrics to quantitatively assess the emotion-evoked image generation performance of all methods. Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions. EMR: We use the emotion predictor $\mathcal{P}$ to predict the emotional category of generated images and calculate the proportion of alignment with the target emotions. ESR: Our emotion predictor $\mathcal{P}$ produces emotion probability distributions $e_{g}$ on generated images and $e_{s}$ on source images. We introduce Kullback-Leibler Divergence distance $KLD(\cdot)$ and calculate ESR as the proportion of the generated images, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$. $e_{oh}$ is the binary one-hot target emotion vector. It indicates the effect of edits introduced on the generated images in transitioning from the source emotion to the target emotion.
Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) measure the structural coherence and semantic consistency between source and generated images. ENRD: We first identify emotionally neutral regions on the source images by employing the Grad-CAM [25] technique through our emotion predictor $\mathcal{P}$. We then binarize the Grad-CAM maps with a threshold of 0.5, where a value of 1 on the resulting binary map signifies regions triggering emotions. By excluding these regions, we compute pixel-level L1 distance between emotionally neutral regions of the source and generated images. ESS: We apply the Canny edge detection algorithm on both source and generated images using thresholds of 200 and 500. Next, we compute the L1 norm between these edge maps to quantify their structural differences.
[Uncaptioned image] Figure 5: The Human Psychophysics Experiment Results. The average proportion of images that human participants prefer our EmoEditor over other competitive methods is 56%. Chance is 50% in the red dotted line.    Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 32.61 7.36 NST 34.42 92.01 34.42 18.57 Csty 11.51 85.52 41.59 36.64 Ip2p 2.53 67.76 9.82 12.71 LMS 11.51 77.38 26.67 19.74 w/o $\mathcal{P}$ 5.06 69.15 20.08 14.93 w/o $L_{emb}$ 41.12 92.36 23.02 16.06 Ours 50.20 92.86 24.73 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor. See Sec. 4 for the methods and evaluation metrics. See Sec. 5.1 for the result analysis.
5 Results
5.1 Quantitative Evaluation in Cross-Valence Scenarios
Figure 5 shows human psychophysics experimental results. If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks. Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec. 4. From Figure 5, we observed that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines. This preference indicates EmoEditor’s excellence in producing images that successfully evoke emotions aligning with the target emotions. This observation is also validated by the highest EMR and ESR scores of our EmoEditor in Table 1. Moreover, our EmoEditor achieves moderately lower ENRD and ESS, suggesting that the generated images by our EmoEditor also preserve structural similarity and semantic coherence.
From Figure 5, we also noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than just global adjustments to color and brightness. Indeed, CT also has lower EMR and ESR than our EmoEditor in Table 1. As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores. Similarly, our EmoEditor is significantly favored over Ip2p by human participants, indicating that current text-to-image diffusion models face challenges in understanding emotional cues and generating effective text instructions for editing emotion-evoked images. Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions.
Additionally, we noted that our EmoEditor also outperforms Large Model Series (LMS) by 4% based on human preference scores in Figure 5, 20% in EMR, 15% in ESR, 2 in ENRD and 3 in ESS in Table 1. It suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
Interestingly, we observed Neural-Style-Transfer [7] and CLIP-Styler [14] excel in inducing negative emotions through distortion and irregular textures on source images. Indeed, they are preferred by humans in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.). However, while these methods create a negative emotional atmosphere, they render the image challenging to comprehend. This is supported by the higher ENRD and ESS scores in Table 1, suggesting that the distorted images no longer preserve the structural similarities and semantic content of source images. In contrast, our EmoEditor outperforms both methods in generating positively valenced images, as supported by higher human preference scores in the lower quadrants of the confusion matrix (see Supp.).
In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance. This leads to lower EMR and ESR scores and higher ENRD and ESS scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions to modify source images, and executing iterative refinements during inference.
Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Sec. 5.2 for the result analysis. See Supp. for more examples.
5.2 Visualization of Emotional Image Generation across Valence
Figure 6 presents visualizations of generated images for all competitive methods in cross-valence scenarios. Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation. For instance, in Row 3, Column 2, it enhances brightness but fails to evoke contentment and amplifies fear-inducing elements in the creepy house. Neural Style Transfer heavily relies on randomly selected reference images. While it produces artistic textures aligned with target emotions, such as Row 1, Column 3, it lacks in preserving essential semantic content in the source images. CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Rows 1-3, Column 4. Ip2p demonstrates limited image generation abilities with single-word emotion prompts. For example, the Halloween scene in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 2, Column 5). Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions. This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 6, Column 6.
In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible. For instance, in Row 1, Column 7, our method ingeniously transforms an unpleasant lunch box into a flower. In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
5.3 Generalization to Real-world Scenarios
Editing images within the same emotional valence presents greater challenges compared to cross-valence emotion image editing due to subtle differences in emotional cues. Figure 7(a) showcases our EmoEditor’s results within the same valence. Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes. In Example 1, it transitions from awe to contentment, depicting a sunset scene with afterglow. In Example 2, it generates a blooming white flower with additional purple flowers to evoke excitement. In Example 3, transitioning from amusement to awe, it transforms a lively spring scene into a tranquil night scene with moonlight.
In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all. We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset [16], commonly used for object detection. These images depict everyday life scenes with neutral emotions. Our EmoEditor applies positive and negative emotion transformations to these source images. Figure 7(b) illustrates the generated results. In Example 1, to inject amusement into an office space, our EmoEditor embellishes the walls with delightful watercolor paintings and infuses lively and vibrant colors into the room furniture. Conversely, to evoke fear, our model creates an aged desk and dimly lit walls. In Example 2, for an image featuring a courtyard with a teddy bear observing the garden, our method enriches the scene by adding a beautiful flower bed, creating an awe-inspiring effect. Conversely, to evoke feelings of disgust, our EmoEditor removes the flowers from the courtyard and transforms the view into a dilapidated scene with shabby clothes hanging on the walls and trash bags in the corner. It is crucial to emphasize that our EmoEditor is not trained in these neutral scenarios. All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios. (a) Three examples are showcased, highlighting our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images. (b) Two examples demonstrate how our EmoEditor can transform neutral real-world images to evoke either positive or negative emotions. Source images are framed in green. Target emotions are in red. See Sec. 5.3 for the result analysis.
6 Discussion
We present a new and important problem of emotion-evoked image generation. To address this, we present EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics. Moreover, we contribute EmoPair dataset for model training. To benchmark all the methods, we introduce new evaluation metrics and establish standard protocols in emotion understanding, visualization, and reasoning. While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images. See Supp. for details.
The problem of emotion-evoked image generation introduced here opens doors for research across AI, psychology, design, arts, and neuroscience. However, EmoEditor, like any technology, carries inherent risks. Misusing it, such as manipulating images to evoke undesirable emotions, could harm individuals’ mental well-being or mislead public feelings and opinions.
Appendix S1 EmoPair Dataset
In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2p[3] and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 [4]. Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten. Figure S1 illustrates examples of these instructions. To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
Figure S1: Instruction examples of EmoPair-Generated Subset (EPGS). For each emotion category, we retained the top ten text instructions based on the rankings determined by human annotators according to the efficacy of each emotion category. We utilized these instructions for Ip2p in image editing to generate target images capable of evoking the desired emotions, thereby constructing our EPGS.
Figure S2: Sample of EmoPair-Annotated Subset (EPAS). On the left side of each pair of images is the source image (framed in green), while the right side shows the target image. The emotion labels for the source and target images (highlighted in red) are indicated above the images.
Figure S3: Sample of EmoPair-Generated Subset (EPGS). On the left side of each pair of images is the source image (framed in green), while the right side shows the target image. The emotion labels for the source and target images (highlighted in red) are indicated above the images.
Appendix S2 Experiment
S2.1 Baseline: Large Model Series
Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS). This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions.
Figure S4 shows the workflow of LMS. First, we employ BLIP to comprehend the source image, generating an image caption corresponding to the source image. Subsequently, based on the generated caption, we utilize the sentence structure “There is an image with [Image caption]. To make me feel [Target Emotion], how to change the image? Give one instruction that starts with an action." to query GPT-3. Following this, GPT-3 generates an instruction to guide Ip2p in editing the source image, resulting in the final output.
Figure S4: Workflow of Large Model Series (LMS). We decompose emotion-evoked image generation into three steps: employing BLIP for image understanding, utilizing GPT-3 for instruction generation, and employing Ip2p for image editing. We refer to this approach of chaining multiple large models as the Large Model Series.
S2.2 Human Psychophysics Experiment
Figure S5(a) shows the MTurk Experiment schematic and Figure S5(b) shows the experiment instruction pages given to the participants. To control the quality of data collected, we have implemented preventive measures to screen participants:
(a) Each participant undergoes 6 randomly dispersed dummy trials within the real experiments. To assess if participants are making random selections, we use 6 image pairs with prominent emotional differences as references. Considering individual emotional response variability, participants are allowed a maximum of one incorrect choice in these dummy trials. Subjects exceeding an error rate of 1/6 are excluded, resulting in 11,400 trials. The outcomes of the 6 dummy trials are also excluded from the final statistics. Figure S5(c) provides an example of a dummy trial.
(b) Each participant can only take part in the experiment once.
(c) Image pairs used in the entire experiment are drawn from the 2,016 result pairs generated by our model and five other SOTA methods. Participants view randomly sampled pairs from these 2,016 pairs, and all trials are presented in a random order. The order of the two images for selection is also randomized. An example trial in the real experiment is illustrated in Figure S5(d).
Figure S5: MTurk Experiment. (a) Schematic of MTurk experiment. Participants are presented with a set of images and a target emotion. They must choose between two images, one generated by our model and the other by five state-of-the-art methods, selecting the one that more effectively evokes the target emotion. (b) Instructions of the MTurk experiment. (c) Dummy trial example. We chose six image pairs with prominent emotional distinctions as benchmarks to assess participants’ comprehension of the task and their attentiveness to the experiment. (d) Real trial example.
Appendix S3 Results
S3.1 Quantitative Evaluation in Cross-Valence Scenarios
We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)[26] (Online). We recruited 417 participants, each undergoing 100 trials. Additionally, we conducted in-lab experiments, involving 10 volunteers. Similarly, each participant reviewed 100 trials. The consistency between online and in-lab experimental results demonstrates the superiority of our approach in producing results that evoke the target emotion compared to other state-of-the-art methods.
Figure S6: Preference of Our Method Over SOTA Methods. We conducted both online experiments (417 participants) and in-lab experiments (10 participants). The left column represents the online results, while the right column displays the in-lab results. The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions. Chance is 50% in the red dotted line.
Figure S6 illustrates the preference of our method over other state-of-the-art methods in both online and in-lab experiments. From the overall results (Row 1), the in-lab experiment demonstrates a higher preference compared to the online experiment. This suggests that our method’s effectiveness in evoking the target emotion is more pronounced in a laboratory setting. Note that our in-lab and online experiments had identical setups. The observed difference may stem from the controlled environment provided by the lab conditions, where participants were more focused, reducing randomness and facilitating a more accurate assessment of emotion generation. This is further supported by the dummy trial success rates, with a 100% pass rate in the in-lab experiment compared to only 27% in the online experiment. Again, we emphasize that we only use the online data for result analysis, where the participants passed the dummy tests with less than 1/6 error rates.
Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-Transfer[7] (NST) and CLIP-Styler[14] (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes. Additionally, our EmoEditor surpasses Color-transfer[21] (CT) and Ip2p[3] in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions. This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones. Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S7 and S8. The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories. From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer [7] and CLIP-Styler [14] more in the manipulation of emotions from positive to negative. This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image. In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images compared to these two approaches.
Figure S7: Confusion matrix of different methods (Online). The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs. The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods. The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods. Average preferences for our EmoEditor and other methods are indicated above each matrix. The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion. Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
Figure S8: Confusion matrix of different methods (In-lab). The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs. The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods. The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods. Average preferences for our EmoEditor and other methods are indicated above each matrix. The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion. Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
S3.2 Visualization of Emotional Image Generation across Valence
Figure S9 illustrates the visualizations of images generated by all methods across various scenarios. Color-Transfer [21] (CT) proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image. For instance, in Row 1, Column 2, it enhances brightness but fails to evoke contentment as it does not alter elements causing disgust, such as the toilet. Neural-Style-Transfer [7] (NST) heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image. For example, in Row 5, Column 3, it generates multiple differently colored blocks, obscuring the original factory background but failing to evoke the emotion of awe. CLIPstyler [14] (Csty) introduces inexplicable textures, as shown in Row 2 and Column 4. Ip2p, due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities, as seen in Column 5. Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion, as evidenced by Row 2, Column 6, where it not only fails to change the crying woman but amplifies the crying details.
In contrast, our EmoEditor, requiring only the target emotion and source image inputs, can generate highly creative images that evoke the target emotion while striving to maintain scene structure and semantic coherence. For example, in Row 1, Column 7, our method creatively transforms the toilet into a table and prepares a delightful breakfast, successfully eliciting a sense of contentment. In Row 2, Column 7, it changes the woman’s pose while retaining the main scene and character from the source image, generating a cartoon-like image that aligns with amusement.
Figure S9: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green.
S3.3 Limitations
Figure S10 shows some failure cases. Our EmoEditor encounters difficulties in handling fine details of individuals, especially when their faces in the images are small. For example, in Row 1, although our EmoEditor attempts to make the boy appear sad, the boy’s face seems somewhat twisted and distorted.
Moreover, significant disparities between source images and target emotions present obstacles, making it difficult to generate results aligning with the target emotion while preserving the structure of the source image. For example, in Row 2, the source image features a tiger, evoking the emotion of anger. In our EmoEditor’s attempt to transform it into awe, although the image’s structure is retained, it is altered to depict steep coastal rocks, changing the semantic content of the image.
Figure S10: Failure Cases. The target emotion is highlighted in red and the source image is framed with green.

### v2

@@
 Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images.
-All code, model, and data will be made public.
+All code, model, and dataset will be made public.
 Figure 1: The generated images evoke a sense of happiness in viewers, contrasting with the negative emotions elicited by the source images.
@@
 For instance, in the first quadrant where trash on the ground evokes negative emotions, our model performs image editing by eliminating the trash while preserving the outdoor context, including trees and grass, in the same spatial arrangement as the original scene.
-For comparisons, we include other competitive methods (Sec.
-4).
+For comparisons, we include other competitive methods (Sec.4).
 Zoom in for better views.
@@
 Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia.
-Despite significant advancements in image generation techniques [6, 27, 35], there has been a notable absence of focus given to emotion conditioning.
+Despite significant advancements in image generation techniques Gal et al.
+(2022); Wang et al.
+(2023); Yu et al.
+(2018), there has been a notable absence of focus given to emotion conditioning.
 Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
@@
 Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response.
-For instance, in Figure 1, the presence of a burning forest significantly contributes to feelings of anger.
-Consequently, in the generated image, the burning forest is substituted with a serene scene featuring a cluster of flowers in the forest, thereby transforming the overall emotional response to awe.
+For instance, in Figure 1, the presence of a burning lawn significantly contributes to feelings of anger.
+Consequently, in the generated image, the burning lawn is substituted with a serene scene featuring a cluster of flowers, thereby transforming the overall emotional response to awe.
+Studies Yang et al.
+(2023, 2018) have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement.
+As also seen in Figure 2, introducing fearful elements like fire or a woman in white into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments.
+However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows.
+In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses.
+Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli Baumeister et al.
+(2001); Ito et al.
+(1998), simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
+Drawing inspiration from these emotion-associated factors, we propose our model, EmoEditor, for emotion-evoked image generation.
+We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties.
+(1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome.
+(2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions.
+(3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
+Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer Pitié et al.
+(2007); Gatys et al.
+(2015); Weng et al.
+(2023), were commonly employed to modify the emotional tone of images.
+These methods typically require reference emotional images or textual cues as additional inputs.
+However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses.
+In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
 Figure 2: Emotions are influenced by both global and local cues from the source images.
@@
 Zoom in for better views.
-Studies [31, 32] have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement.
-As also seen in Figure 2, introducing nauseating or fearful elements like fire or ghosts into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments.
-However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows.
-In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses.
-Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli [2, 11], simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
-Drawing inspiration from these emotion-associated factors above, we propose our model, EmoEditor, for emotion-evoked image generation.
-We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties.
-(1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome.
-(2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions for image editing.
-(3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
-Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer [21, 7, 30], were commonly employed to modify the emotional tone of images.
-These methods typically require reference emotional images or textual cues as additional inputs.
-However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses.
-In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
-With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged [12, 3, 24].
+With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged Kawar et al.
+(2023); Brooks et al.
+(2023); Ruiz et al.
+(2023).
 However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images.
@@
 To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions.
-Unlike existing datasets on visual emotion analysis [18, 17, 34, 20, 31], which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
+Unlike existing datasets on visual emotion analysis Mikels et al.
+(2005); Machajdik, Hanbury (2010); You et al.
+(2016); Peng et al.
+(2015); Yang et al.
+(2023), which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
 The main contributions of this paper are highlighted below: (1) We introduce a new and important problem of emotion-evoked image generation.
@@
 2.1 Visual Emotion Analysis
-Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses through visual information from images or videos.
-In psychology, Categorical Emotion States (CES) form the basis for emotion classification models, notably exemplified by the 6-category Ekman model [5] and the 8-category Mikels model [18].
-Early research focused on manually crafted features such as color and content [17], while recent advancements leverage deep learning models to automatically capture intrinsic global and local emotional cues for more accurate predictions [33, 32].
-Despite advancements in visual emotion analysis, there is a limited focus on generating images that evoke specific emotional responses.
-To address this, we propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation in a unified framework.
-Unlike Affective Image Filter (AIF) [30] which relies on explicit textual instructions, EmoEditor does not need explicit text guidance on scene modifications for desired emotions.
-Exsiting VEA datasets [20, 17, 34] primarily focus on emotion classification, including EmoSet [31], the largest dataset with 120,000 images annotated for emotion labels and six attributes.
-However, none of them include image pairs with annotated emotion labels and text editing instructions crucial for emotion-evoked image generation.
-Therefore, we introduce the EmoPair dataset, comprising 340,000 image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
+Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses from visual data.
+In psychology, Categorical Emotion States (CES) form the basis for emotion classification models like the 8-category Mikels model Mikels et al.
+(2005).
+Early research used manually crafted features such as color and content Machajdik, Hanbury (2010), while recent advancements employ deep learning to capture emotional cues more accurately You et al.
+(2015); Yang et al.
+(2018).
+Despite progress, generating images to evoke specific emotions remains underexplored.
+We propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation to address this.
+Unlike the Affective Image Filter Weng et al.
+(2023), which requires explicit textual instructions, EmoEditor does not need text guidance to achieve desired emotions.
+Existing VEA datasets Peng et al.
+(2015); You et al.
+(2016) focus on emotion classification, including EmoSet Yang et al.
+(2023), which has 120k images with emotion labels and attributes.
+However, they lack image pairs with annotated emotions and text editing instructions essential for emotion-evoked image generation.
+Thus, we introduce the EmoPair dataset, comprising 340k image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
 2.2 Image Editing
 Color and Style Transfer.
-Color and style transfer methods like PDF-Transfer[21], Neural Style Transfer[7], and CLIP-Styler[14] use reference images for image manipulation, with outcomes varying based on choices.
-While colors and styles may influence human emotions, they struggle to manipulate local image regions, limiting their versatility in emotion-evoked generation.
-In contrast, our EmoEditor balances local and global image manipulations during inference without requiring additional references.
+Methods like PDF-TransferPitié et al.
+(2007), Neural Style TransferGatys et al.
+(2015), and CLIP-StylerKwon, Ye (2022) manipulate images using reference images, but they struggle with local region adjustments, limiting their versatility in generating emotion-evoking images.
+Our EmoEditor balances local and global manipulations without needing additional references.
 Text-to-Image Editing.
-Text-to-image editing is a growing area of interest.
-Previous studies utilize Generative Adversarial Networks (GAN) [8] architectures and leverage CLIP model’s [22] language embedding for image manipulation from detailed text descriptions[19, 1].
-Recently, text-to-image diffusion models like Imagic[12], Prompt-to-prompt[10], InstructPix2Pix[3], and DreamBooth[24] have strong capabilities in generating high-quality and diverse images based on text prompts.
-Despite their impressive results, they often overlook emotional aspects and rely heavily on detailed textual instructions.
-In contrast, our EmoEditor only needs target emotion and source image inputs.
-Through end-to-end training, it comprehends emotional cues and creatively edits images for the desired emotions while preserving structure and coherence.
+Text-to-image editing is increasingly popular.
+Earlier studies used GANs Goodfellow et al.
+(2014) and CLIP models Radford et al.
+(2021) for image manipulation based on detailed text descriptions Patashnik et al.
+(2021); Abdal et al.
+(2022).
+Recent diffusion models like Imagic Kawar et al.
+(2023), Prompt-to-prompt Hertz et al.
+(2022), InstructPix2Pix Brooks et al.
+(2023), and DreamBooth Ruiz et al.
+(2023) generate high-quality images from text prompts but often ignore emotional aspects and rely heavily on detailed instructions.
+In contrast, our EmoEditor needs only target emotion and source image inputs, using end-to-end training to understand and edit emotional cues while preserving image structure and coherence.
 3 Method
 We formulate emotion-evoked image generation as a supervised learning task.
-First, we curate the EmoPair dataset containing image pairs with emotion labels and corresponding text instructions for editing (Sec.
-3.1).
-Next, EmoEditor is trained on this dataset (Sec.
-3.2).
+First, we curate the EmoPair dataset containing image pairs with emotion labels and corresponding text instructions for editing (Sec.3.1).
+Next, EmoEditor is trained on this dataset (Sec.3.2).
 During inference, it takes source images and target emotions, producing generated images without the need for human-annotated text instructions or emotion reference images.
@@
 The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided.
-See more examples in Supp..
-See Sec.
-3.1 for EmoPair dataset.
+See more examples in Supp.Sec.S1.
+See Sec.3.1 for EmoPair dataset.
 3.1 Curating Our EmoPair Dataset
 Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair.
-It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2p[3] annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions.
+It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2pBrooks et al.
+(2023) annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions.
 All images from EmoPair are of size 224$\times$224.
-Similar to existing works in VEA [18, 31], all images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions.
 Figure 3 shows the pipeline for curating our EmoPair dataset.
+Emotion Predictor $\mathcal{P}$.
+To better construct our dataset, we first train an emotion predictor $\mathcal{P}$ on the existing VEA dataset EmoSet Yang et al.
+(2023).
+All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions.
+$\mathcal{P}$ employs ResNet18 He et al.
+(2016) as a backbone and achieves a top-1 accuracy of 73% on EmoSet’s test set.
 EmoPair-Annotated Subset (EPAS).
-The original Ip2p dataset [3] lacks emotion labels, consisting only of image pairs with text instructions for editing.
-To address this limitation, we use an emotion predictor $\mathcal{P}$ to classify the source and target images from the Ip2p dataset into eight distinct emotions.
-$\mathcal{P}$ employs ResNet18 [9] as a backbone and trained on the EmoSet dataset [31].
-It achieves a top-1 accuracy of 73% on EmoSet’s test set.
-We augment the Ip2p dataset with emotion labels, resulting in a total of 331,595 image pairs.
-Notably, the text instructions with these image pairs serve as human expectations to guide EmoEditor’s training (Sec.
-3.2), but are not utilized during inference.
+The original Ip2p dataset lacks emotion labels, consisting only of image pairs with text instructions for editing.
+To address this, we use $\mathcal{P}$ to classify the source and target images from the Ip2p dataset into eight distinct emotions.
+We augment the Ip2p dataset with emotion labels, resulting in 331,595 image pairs.
+Notably, the text instructions with these image pairs serve as human expectations to guide EmoEditor’s training but are not utilized during inference.
 EmoPair-Generated Subset (EPGS).
-Due to the lack of salient emotional cues in the Ip2p dataset, we further complement EPAS with EPGS subset, which has emotionally rich images from the EmoSet dataset [31].
-EmoSet provides diverse and evocative emotional cues that benefit our EmoEditor.
-However, EmoSet lacks paired images with coherent scene structures and target emotions, along with the absence of text editing instructions from source to target images.
-Directly inputting a single-word prompt like “[Target Emotion]" into existing text-to-image editing models proves ineffective due to their limited emotional understanding and reasoning capabilities.
+To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet Yang et al.
+(2023).
+However, EmoSet lacks paired images with consistent scenes and target emotions, and it doesn’t include text editing instructions.
+Single-word prompts like “[Target Emotion]" prove ineffective in existing text-to-image models due to their limited emotional reasoning.
 Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images.
 For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape.
-Thus, to ensure accurate image pair generation, we created 50 general instructions, agnostic of source images, to prompt transitions to desired emotions across 8 categories using GPT-3 [4].
-Human annotators then ranked these instructions per emotion category based on efficacy and retained the top ten.
-Next, the pre-trained Ip2p model manipulated EmoSet source images using these selected instructions to generate target images.
-We employ the following selection criteria to control the quality of generated image pairs: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions.
-(2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) [29] and Learned Perceptual Image Patch Similarity (LPIPS) [36] for filtering the remaining outcomes.
-SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences.
-Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1.
-Ultimately, the EPGS subset retains 6,949 image pairs.
+Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 Brown et al.
+(2020).
+Human annotators rank these instructions, keeping the top ten per emotion.
+The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures (see Supp.Sec.S1).
 3.2 Our Diffusion Model - EmoEditor
@@
 Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion.
-It employs the latent diffusion model (LDM) [23] with three key novelties.
+It employs the latent diffusion model (LDM) Rombach et al.
+(2022) with three key novelties.
 EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios.
@@
 In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0.
-Next, we introduce an emotion encoder $\tau_{\theta}$, which encodes $e_{oh}$ to $e$ via a stack of fully connected blocks: $e=\tau_{\theta}(e_{oh})$.
-Unlike classical text-to-image models using explicit text prompts, our $\tau_{\theta}$ is trained from scratch, learning neuro-symbolic embeddings for the reverse diffusion process without relying on pre-trained language models like CLIP[22].
+This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation.
+Conversely, text inputs demand extra transformation steps, potentially introducing noise or errors.
+We introduce an emotion encoder $\tau_{\theta}$ to encode $e_{oh}$ into $e$ via fully connected layers: $e=\tau_{\theta}(e_{oh})$ (see Supp.Sec.S2.1 for details), trained from scratch to learn neuro-symbolic embeddings.
 Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$.
-Notably, we intentionally avoid using explicit text editing instructions from image pairs as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions.
+Notably, we intentionally avoid using explicit text editing instructions as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions.
 Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
 In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$.
-Similar to the LDM [23], $\mathcal{E}$ is based on a pre-trained variational autoencoder[13].
-The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps and $t\in T$.
-Specifically, we expressed $z_{t}$ as $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution with mean 0 and variance 1.
+Similar to the LDM Rombach et al.
+(2022), $\mathcal{E}$ is based on a pre-trained VAEKingma, Welling (2022).
+The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps ($t\in T$).
+Specifically, we expressed $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution.
 In the reverse diffusion process, the denoising network $\epsilon_{\theta}$ predicts noise added to the noisy latent $z_{t}$ using cross-attention with the target emotion condition $e$ and source image condition $z$.
@@
 In contrast to existing text-to-image models, EmoEditor emulates human reasoning by aligning its thought process with text instructions in our EmoPair dataset, which leads the transition from the source emotion to the target emotion.
-To achieve this, we introduce an alignment loss in a neuro-symbolic space, between emotion embeddings $e$ from $\tau_{\theta}$ and text embeddings $c$ of instructions in each image pair of our EmoPair dataset with pre-trained CLIP [22].
+To achieve this, we introduce an alignment loss in a neuro-symbolic space, between emotion embeddings $e$ from $\tau_{\theta}$ and text embeddings $c$ of instructions in each image pair of our EmoPair dataset with pre-trained CLIP.
 Specifically, the alignment loss is formulated as the inverse of cosine similarity between $e$ and $c$: $L_{emb}=1-\cos(e,c)$.
@@
 Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic.
-This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 20.
+This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 30.
 In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
-Implementation Details.
-Follow Wang et al.[28], we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p[3] weights.
-Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$.
-The frozen emotion predictor $\mathcal{P}$ is only used during inference.
-We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer.
-The max time step $T$ is set to 1,000.
+See Supp.Sec.S2.2 for Implementation Details.
 4 Experiment
-We selected 504 images randomly from the EmoSet dataset [31], ensuring they were distinct from those in our EmoPair dataset.
+We selected 504 images randomly from the EmoSet dataset Yang et al.
+(2023), ensuring they were distinct from those in our EmoPair dataset.
 According to the emotion labels of EmoSet, there are 63 images for each source emotion category.
@@
 Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween.
-Secondly, images of the same valence may contain subtle differences challenging for human viewers to discern.
+Secondly, images of the same valence may contain subtle differences challenging for humans to discern.
 For instance, an amusement park scene might evoke a mixture of excitement and amusement.
 Baselines.
-We compare our EmoEditor with five state-of-the-art methods: (1) Color-transfer[21] (CT); (2) Neural-Style-Transfer[7] (NST).
+We compare our EmoEditor with five state-of-the-art methods: (1) Color-transferPitié et al.
+(2007) (CT); (2) Neural-Style-TransferGatys et al.
+(2015) (NST).
 Both methods require an additional reference image from the target emotion category as inputs.
 We randomly selected these reference images from the EmoSet dataset.
-We also include representative text-to-image editing models (3) CLIP-Styler[14] (Csty) and (4) Ip2p [3].
+We also include representative text-to-image editing models (3) CLIP-StylerKwon, Ye (2022) (Csty) and (4) Ip2p Brooks et al.
+(2023).
 We provide these models with the target emotion category as a one-word text prompt, such as “excitement".
@@
 To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS).
-This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions.
-See Supp.
-for details.
+This includes BLIP Li et al.
+(2022) for image captioning, followed by GPT-3 Brown et al.
+(2020) for text instruction generation, and Ip2p for image editing based on the instructions (See Supp.Sec.S2.3).
+[Uncaptioned image] Figure 5: Human Psychophysics Experiment Results.
+The average proportion of images that human participants prefer our EmoEditor over other methods is 56%.
+Chance is 50% (red dotted line).
+Error bars are standard errors.
+Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 33.29 7.36 NST 34.42 92.01 34.45 18.57 Csty 11.51 85.52 41.47 36.64 Ip2p 2.53 67.76 9.39 12.71 LMS 11.51 77.38 26.13 19.74 w/o $\mathcal{P}$ 5.06 69.15 19.45 14.93 w/o $L_{emb}$ 43.35 91.62 22.53 16.00 Ours 50.20 92.86 23.98 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor.
+See Sec.4 for the methods and evaluation metrics.
+See Sec.5.1 for the result analysis.
 Human Psychophysics Experiment.
-We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)[26].
+We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)Turk (2012).
 We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials.
@@
 We discard the data from those people who fail the control trials at least 17% of the time.
-See Supp.
-for more experimental details.
+See Supp.Sec.S2.4 for more details.
 Evaluation Metrics.
-We introduce four novel evaluation metrics to quantitatively assess the emotion-evoked image generation performance of all methods.
+While CLIP Radford et al.
+(2021) is commonly used to assess text-image consistency, it doesn’t fully capture the complexity of emotions Widhoelzl, Takmaz (2024).
+Relying on CLIP for emotional alignment may overlook important nuances in images.
+Thus, we introduce four novel evaluation metrics to quantitatively assess emotion-evoked image generation performance.
 Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions.
-EMR: We use the emotion predictor $\mathcal{P}$ to predict the emotional category of generated images and calculate the proportion of alignment with the target emotions.
-ESR: Our emotion predictor $\mathcal{P}$ produces emotion probability distributions $e_{g}$ on generated images and $e_{s}$ on source images.
-We introduce Kullback-Leibler Divergence distance $KLD(\cdot)$ and calculate ESR as the proportion of the generated images, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$.
-$e_{oh}$ is the binary one-hot target emotion vector.
-It indicates the effect of edits introduced on the generated images in transitioning from the source emotion to the target emotion.
-Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) measure the structural coherence and semantic consistency between source and generated images.
-ENRD: We first identify emotionally neutral regions on the source images by employing the Grad-CAM [25] technique through our emotion predictor $\mathcal{P}$.
-We then binarize the Grad-CAM maps with a threshold of 0.5, where a value of 1 on the resulting binary map signifies regions triggering emotions.
-By excluding these regions, we compute pixel-level L1 distance between emotionally neutral regions of the source and generated images.
-ESS: We apply the Canny edge detection algorithm on both source and generated images using thresholds of 200 and 500.
-Next, we compute the L1 norm between these edge maps to quantify their structural differences.
-[Uncaptioned image] Figure 5: The Human Psychophysics Experiment Results.
-The average proportion of images that human participants prefer our EmoEditor over other competitive methods is 56%.
-Chance is 50% in the red dotted line.
-Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 32.61 7.36 NST 34.42 92.01 34.42 18.57 Csty 11.51 85.52 41.59 36.64 Ip2p 2.53 67.76 9.82 12.71 LMS 11.51 77.38 26.67 19.74 w/o $\mathcal{P}$ 5.06 69.15 20.08 14.93 w/o $L_{emb}$ 41.12 92.36 23.02 16.06 Ours 50.20 92.86 24.73 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor.
-See Sec.
-4 for the methods and evaluation metrics.
-See Sec.
-5.1 for the result analysis.
+EMR: Similar to Goetschalckx et al.
+(2019), we also use our emotion predictor $\mathcal{P}$ to assess the emotional category of generated images and determine their alignment with target emotions.
+We acknowledge that using $\mathcal{P}$ as an evaluator may introduce favorable biases toward our model.
+However, the absence of existing emotion evaluators leaves us no choice but to rely on $\mathcal{P}$.
+ESR: $\mathcal{P}$ generates emotion probability distributions $e_{g}$ and $e_{s}$ for generated and source images.
+We compute ESR using Kullback-Leibler Divergence distance $KLD(\cdot)$, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$.
+$e_{oh}$ is the binary one-hot target emotion vector, indicating the impact of edits on generated images transitioning from the source emotion to the target emotion.
+Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) assess the structural coherence and semantic consistency between source and generated images.
+ENRD: Using Grad-CAM Selvaraju et al.
+(2017) with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify emotionally neutral regions on source images, valued at 0.
+Then, we compute the pixel-level L1 distance between these regions on source and generated images.
+ESS: Employing the Canny edge detection algorithm with thresholds of 200 and 500 on both source and generated images, we determine the L1 norm between their edge maps to quantify structural disparities.
 5 Results
 5.1 Quantitative Evaluation in Cross-Valence Scenarios
-Figure 5 shows human psychophysics experimental results.
+Figure 5 shows human psychophysics experimental results (see Supp.Sec.S3.1 for more results).
 If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks.
-Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec.
-4.
-From Figure 5, we observed that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines.
-This preference indicates EmoEditor’s excellence in producing images that successfully evoke emotions aligning with the target emotions.
-This observation is also validated by the highest EMR and ESR scores of our EmoEditor in Table 1.
-Moreover, our EmoEditor achieves moderately lower ENRD and ESS, suggesting that the generated images by our EmoEditor also preserve structural similarity and semantic coherence.
-From Figure 5, we also noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than just global adjustments to color and brightness.
-Indeed, CT also has lower EMR and ESR than our EmoEditor in Table 1.
+Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec.4.
+Observing Figure 5, we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines, indicating its proficiency in evoking target emotions.
+This is supported by EmoEditor’s highest EMR and ESR scores in Table 1.
+Additionally, EmoEditor achieves moderately lower ENRD and ESS, suggesting maintained structural and semantic coherence in the generated images.
+Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Supp.Fig.S10 for more examples.
+From Figure 5, we noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than global color and brightness adjustments.
+CT also has lower EMR and ESR than our EmoEditor in Table 1.
 As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores.
@@
 Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions.
-Additionally, we noted that our EmoEditor also outperforms Large Model Series (LMS) by 4% based on human preference scores in Figure 5, 20% in EMR, 15% in ESR, 2 in ENRD and 3 in ESS in Table 1.
-It suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
-Interestingly, we observed Neural-Style-Transfer [7] and CLIP-Styler [14] excel in inducing negative emotions through distortion and irregular textures on source images.
-Indeed, they are preferred by humans in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.).
-However, while these methods create a negative emotional atmosphere, they render the image challenging to comprehend.
-This is supported by the higher ENRD and ESS scores in Table 1, suggesting that the distorted images no longer preserve the structural similarities and semantic content of source images.
-In contrast, our EmoEditor outperforms both methods in generating positively valenced images, as supported by higher human preference scores in the lower quadrants of the confusion matrix (see Supp.).
+Additionally, our EmoEditor also outperforms Large Model Series (LMS) based on human preference scores in Figure 5, 39% in EMR, 15% in ESR, 2 in ENRD, and 3 in ESS in Table 1.
+This suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
+Neural-Style-Transfer (NST) and CLIP-Styler (Csty) excel in inducing negative emotions through distortion and irregular textures on source images.
+They are preferred in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.Fig.S8).
+However, these methods render images challenging to comprehend, indicated by higher ENRD and ESS scores in Table 1, suggesting the distorted images lose structural similarities and semantic content.
+In contrast, our EmoEditor outperforms both methods in generating positively valenced images, with higher preference scores in the lower quadrants of the confusion matrix (see Supp.Fig.S8).
 In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance.
-This leads to lower EMR and ESR scores and higher ENRD and ESS scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions to modify source images, and executing iterative refinements during inference.
-Figure 6: Visualization of Generated Images from Different Methods.
-The target emotion is highlighted in red and the source image is framed with green.
-See Sec.
-5.2 for the result analysis.
-See Supp.
-for more examples.
+This leads to lower EMR and ESR scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions, and executing iterative refinements during inference.
+We also perform ablation studies on different max iterations for iterative emotion inference, showing that this iterative process improves the model’s performance, achieving stable and high-quality emotionally evocative images (see Supp.Sec.S3.3).
 5.2 Visualization of Emotional Image Generation across Valence
-Figure 6 presents visualizations of generated images for all competitive methods in cross-valence scenarios.
+Figure 1 and 6 present visualizations of generated images for all competitive methods in cross-valence scenarios (see Supp.Sec.S3.2 for more results).
 Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation.
-For instance, in Row 3, Column 2, it enhances brightness but fails to evoke contentment and amplifies fear-inducing elements in the creepy house.
 Neural Style Transfer heavily relies on randomly selected reference images.
-While it produces artistic textures aligned with target emotions, such as Row 1, Column 3, it lacks in preserving essential semantic content in the source images.
-CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Rows 1-3, Column 4.
+While it produces artistic textures aligned with target emotions, it lacks in preserving essential semantic content in the source images.
+CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Column 4.
 Ip2p demonstrates limited image generation abilities with single-word emotion prompts.
-For example, the Halloween scene in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 2, Column 5).
+For example, the fire in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 1, Column 5).
 Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions.
-This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 6, Column 6.
+This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 2, Column 6.
 In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible.
-For instance, in Row 1, Column 7, our method ingeniously transforms an unpleasant lunch box into a flower.
+For instance, in Row 1, Column 7, our method does not change the subject of the fire but turns the fire in the woods into a bonfire and generates a beautiful night sky in the background, thus evoking contentment.
 In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
+Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios.
+(a) Three examples are showcased, highlighting our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images.
+(b) Two examples demonstrate how our EmoEditor can transform neutral real-world images to evoke either positive or negative emotions.
+Source images are framed in green.
+Target emotions are in red.
+See Sec.5.3 for the result analysis.
 5.3 Generalization to Real-world Scenarios
@@
 Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes.
-In Example 1, it transitions from awe to contentment, depicting a sunset scene with afterglow.
-In Example 2, it generates a blooming white flower with additional purple flowers to evoke excitement.
-In Example 3, transitioning from amusement to awe, it transforms a lively spring scene into a tranquil night scene with moonlight.
+Example 1 transitions from awe to contentment, depicting a sunset scene with afterglow.
+Example 2 generates a blooming white flower with additional purple flowers to evoke excitement.
+Example 3 transforms a lively spring scene into a tranquil night scene with moonlight to evoke awe.
 In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all.
-We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset [16], commonly used for object detection.
+We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset Lin et al.
+(2014), commonly used for object detection.
 These images depict everyday life scenes with neutral emotions.
@@
 All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
-Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios.
-(a) Three examples are showcased, highlighting our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images.
-(b) Two examples demonstrate how our EmoEditor can transform neutral real-world images to evoke either positive or negative emotions.
-Source images are framed in green.
-Target emotions are in red.
-See Sec.
-5.3 for the result analysis.
 6 Discussion
@@
 While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images.
-See Supp.
-for details.
+See Supp.Sec.S3.4 for details.
 The problem of emotion-evoked image generation introduced here opens doors for research across AI, psychology, design, arts, and neuroscience.
@@
 Misusing it, such as manipulating images to evoke undesirable emotions, could harm individuals’ mental well-being or mislead public feelings and opinions.
+Supplementary Material
 Appendix S1 EmoPair Dataset
-In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2p[3] and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
-For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 [4].
+In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2pBrooks et al.
+(2023) and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
+For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 Brown et al.
+(2020).
 Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten.
 Figure S1 illustrates examples of these instructions.
+We employ the following selection criteria to control the quality of generated image pairs of EPGS: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions.
+(2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) Wang et al.
+(2004) and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al.
+(2018) for filtering the remaining outcomes.
+SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences.
+Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1.
+Ultimately, EPGS retains 6,949 image pairs.
 To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
@@
 Appendix S2 Experiment
-S2.1 Baseline: Large Model Series
+S2.1 Emotion Encoder
+Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming one-hot encoded emotion input vectors into the structured emotion embedding $e$.
+The architecture of the model is composed of a series of fully connected layers that progressively increase the dimensionality of the input.
+Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$.
+The network begins with an input layer of size 8, which is then passed through a sequence of fully connected layers with increasing sizes: 256, 512, and 768 neurons, respectively.
+Each of these layers is followed by a ReLU activation function to introduce non-linearity and improve the model’s learning capacity.
+The final linear layer projects the output to a dimension of $77\times 768$, effectively structuring the embedding to fit the input size of our denoising network $\epsilon_{\theta}$.
+Figure S4: The network structure of our emotion encoder $\tau_{\theta}$.
+S2.2 Implementation Details
+Follow Wang et al.
+(2022), we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p Brooks et al.
+(2023) weights.
+Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$.
+The frozen emotion predictor $\mathcal{P}$ is only used during inference.
+We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer.
+The max time step $T$ is set to 1,000.
+S2.3 Baseline: Large Model Series
 Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing.
 To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS).
-This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions.
+This includes BLIP Li et al.
+(2022) for image captioning, followed by GPT-3 Brown et al.
+(2020) for text instruction generation, and Ip2p for image editing based on the instructions.
 Figure S5 shows the workflow of LMS.
@@
 We refer to this approach of chaining multiple large models as the Large Model Series.
-S2.2 Human Psychophysics Experiment
+S2.4 Human Psychophysics Experiment
 Figure S6(a) shows the MTurk Experiment schematic and Figure S6(b) shows the experiment instruction pages given to the participants.
@@
 S3.1 Quantitative Evaluation in Cross-Valence Scenarios
-We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)[26] (Online).
+We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)Turk (2012) (Online).
 We recruited 417 participants, each undergoing 100 trials.
@@
 Chance is 50% in the red dotted line.
+Error bars are standard errors.
 Figure S7 illustrates the preference of our method over other state-of-the-art methods in both online and in-lab experiments.
@@
 Again, we emphasize that we only use the online data for result analysis, where the participants passed the dummy tests with less than 1/6 error rates.
-Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-Transfer[7] (NST) and CLIP-Styler[14] (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes.
-Additionally, our EmoEditor surpasses Color-transfer[21] (CT) and Ip2p[3] in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions.
+Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-TransferGatys et al.
+(2015) (NST) and CLIP-StylerKwon, Ye (2022) (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes.
+Additionally, our EmoEditor surpasses Color-transferPitié et al.
+(2007) (CT) and Ip2pBrooks et al.
+(2023) in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions.
 This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones.
 Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
-In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S7 and S8.
+In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S8 and S9.
 The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories.
-From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer [7] and CLIP-Styler [14] more in the manipulation of emotions from positive to negative.
+From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer Gatys et al.
+(2015) and CLIP-Styler Kwon, Ye (2022) more in the manipulation of emotions from positive to negative.
 This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image.
@@
 Figure S10 illustrates the visualizations of images generated by all methods across various scenarios.
-Color-Transfer [21] (CT) proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image.
-For instance, in Row 1, Column 2, it enhances brightness but fails to evoke contentment as it does not alter elements causing disgust, such as the toilet.
-Neural-Style-Transfer [7] (NST) heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image.
-For example, in Row 5, Column 3, it generates multiple differently colored blocks, obscuring the original factory background but failing to evoke the emotion of awe.
-CLIPstyler [14] (Csty) introduces inexplicable textures, as shown in Row 2 and Column 4.
-Ip2p, due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities, as seen in Column 5.
-Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion, as evidenced by Row 2, Column 6, where it not only fails to change the crying woman but amplifies the crying details.
+Color-Transfer Pitié et al.
+(2007) (CT) proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image.
+For instance, in Row 1, Column 2, it enhances brightness but fails to evoke amusement as it does not alter elements causing sadness, such as the crying woman.
+Neural-Style-Transfer Gatys et al.
+(2015) (NST) heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image.
+For example, in Row 4, Column 3, it generates multiple differently colored blocks but fails to evoke the emotion of sadness.
+CLIPstyler Kwon, Ye (2022) (Csty) introduces inexplicable textures.
+Ip2p Brooks et al.
+(2023), due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities, as seen in Column 5.
+Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion, as evidenced by Row 1, Column 6, where it not only fails to change the crying woman but amplifies the crying details.
+In addition, it often greatly changes the scene of the source image to achieve the goal of changing the emotion, such as Row 2, Column 6, it produces a building that looks very technological but quite different from the source image.
 In contrast, our EmoEditor, requiring only the target emotion and source image inputs, can generate highly creative images that evoke the target emotion while striving to maintain scene structure and semantic coherence.
-For example, in Row 1, Column 7, our method creatively transforms the toilet into a table and prepares a delightful breakfast, successfully eliciting a sense of contentment.
-In Row 2, Column 7, it changes the woman’s pose while retaining the main scene and character from the source image, generating a cartoon-like image that aligns with amusement.
+For example, in Row 1, Column 7, it changes the woman’s pose while retaining the main scene and character from the source image, generating a cartoon-like image that aligns with amusement.
+In Row 2, Column 7, our EmoEditor generates technological results that make people feel awe while almost maintaining most of the structure of the source image.
 Figure S10: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green.
-S3.3 Limitations
+S3.3 Ablation Study of Different Iterations
+Due to the abstract and complex nature of emotions, we propose an iterative emotion inference approach.
+The maximum number of iterations significantly impacts the quality of images generated by the model.
+To assess this effect, we vary the maximum number of iterations from 1 to 30 and evaluate the results using four metrics.
+Figure S11 presents our experimental findings.
+The Emotion Matching Ratio (EMR) shows a steady increase from 0% to approximately 50% over 30 iterations, indicating continuous improvement in the model’s ability to generate images that evoke the target emotions.
+The Emotion Similarity Ratio (ESR) rises rapidly in the initial iterations, stabilizing around 90%.
+This suggests that the model quickly becomes proficient at producing images that closely match the target emotional content, with minimal improvements after the initial iterations.
+The Emotion-Neutral Region Deviation (ENRD) also increases rapidly at first, plateauing around 22.5.
+This pattern indicates that the deviation in neutral regions of the generated images becomes consistent after a few iterations, reflecting the model’s stability in preserving non-emotional aspects.
+The Edge Structure Similarity (ESS) follows a similar trend, rising quickly to around 16 and then stabilizing.
+This stability in ESS reflects the model’s ability to maintain structural coherence and semantic consistency between the source and generated images after the initial iterations.
+In summary, these metrics demonstrate that the iterative process enhances the model’s performance significantly in the initial stages, reaching a stable state of high quality and consistency in generating emotionally evocative images.
+Figure S11: Performance Metrics Across Different Iterations.
+S3.4 Limitations
 Figure S12 shows some failure cases.
@@
 Figure S12: Failure Cases. The target emotion is highlighted in red and the source image is framed with green.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The main claims made in the abstract and introduction accurately align with the paper’s contributions and scope, reflecting both experimental results as well as acknowledging important assumptions and limitations.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We mentioned the limitations of the paper in Sec.6 of the main text and in Sec.S3.4 of the supplementary materials.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We provided a clear description of our model and dataset in Sec.3, and elaborated on our experimental details in the supplementary materials Sec.S2.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: All code, model, and dataset will be made public.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We detailed the construction of the dataset, the model’s architecture, and the training and inference methods in Sec.3, with further elaboration provided in Sec.S2 of the supplementary materials.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We reported error bars for the experiments in Figure 5 and Figure S7, using standard error.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Implementation details are shown in Sec.S2.2.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We mentioned the broader impacts of the paper in Sec.6.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [Yes]
+  54. Justification: The model will require users to adhere to usage guidelines upon its release.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All utilized datasets are properly cited in the paper and are solely used for academic research purposes.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: All code, model, and dataset will be made public.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [Yes]
+  69.
+Justification: All detailed information regarding human psychophysics experiments is presented in Sec.S2.4, including experimental setup, procedures, interface design, and results.
+All participants are properly compensated.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [Yes]
+  74. Justification: All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 Figure S12: Failure Cases. The target emotion is highlighted in red and the source image is framed with green.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The main claims made in the abstract and introduction accurately align with the paper’s contributions and scope, reflecting both experimental results as well as acknowledging important assumptions and limitations.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We mentioned the limitations of the paper in Sec.6 of the main text and in Sec.S3.4 of the supplementary materials.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We provided a clear description of our model and dataset in Sec.3, and elaborated on our experimental details in the supplementary materials Sec.S2.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: All code, model, and dataset will be made public.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: We detailed the construction of the dataset, the model’s architecture, and the training and inference methods in Sec.3, with further elaboration provided in Sec.S2 of the supplementary materials.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We reported error bars for the experiments in Figure 5 and Figure S7, using standard error.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Implementation details are shown in Sec.S2.2.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We mentioned the broader impacts of the paper in Sec.6.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: The model will require users to adhere to usage guidelines upon its release.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All utilized datasets are properly cited in the paper and are solely used for academic research purposes.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: All code, model, and dataset will be made public.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [Yes]
-  69.
-Justification: All detailed information regarding human psychophysics experiments is presented in Sec.S2.4, including experimental setup, procedures, interface design, and results.
-All participants are properly compensated.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [Yes]
-  74. Justification: All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v4

@@
 The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design.
-For the first time, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes.
+First, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes.
 To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments.
 Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations.
-Furthermore, we conduct human psychophysics experiments and introduce four new evaluation metrics to systematically benchmark all the methods.
+Furthermore, we conduct human psychophysics experiments and introduce a new evaluation metric to systematically benchmark all the methods.
 Experimental results demonstrate that our method surpasses all competitive baselines.
 Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images.
-All code, model, and dataset will be made public.
+All code, model, and dataset are available at GitHub.
 Figure 1: The generated images evoke a sense of happiness in viewers, contrasting with the negative emotions elicited by the source images.
 Given a source image that triggers negative emotions (framed in green), our method (Ours) synthesizes a new image that elicits the given positive target emotions (in red), while maintaining the essential elements and structures of the scene.
-For instance, in the first quadrant where trash on the ground evokes negative emotions, our model performs image editing by eliminating the trash while preserving the outdoor context, including trees and grass, in the same spatial arrangement as the original scene.
-For comparisons, we include other competitive methods (Sec.4).
-Zoom in for better views.
+For instance, our method replaces the anger-inducing flames in the source image with cute-shaped lamps to evoke the target emotion of amusement.
+While in an outdoor setting, the raging fire is substituted with a tranquil, lush meadow to inspire a sense of awe.
+For comparisons, we include other competitive methods.
+The blue number below each image represents its CAM-based ESMI score, with higher values being better.
 1 Introduction
-“I am feeling down. Enhance my room to bring in more excitement."
+“I am feeling down. Bring some excitement to my room."
 What we perceive not only entails useful visual information but also evokes profound emotional reactions.
 Incorporating visual cues that elicit emotions into images holds practical significance in various domains.
-For instance, in commercials, it amplifies product branding and shapes public opinions, captivating consumer attention.
-Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia.
-Despite significant advancements in image generation techniques Gal et al.
-(2022); Wang et al.
-(2023); Yu et al.
-(2018), there has been a notable absence of focus given to emotion conditioning.
-Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
-Emotion-evoked image generation poses a challenge as it requires a deep understanding and identification of the subtle context and semantic elements present in the source images that evoke emotional responses.
-Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response.
-For instance, in Figure 1, the presence of a burning lawn significantly contributes to feelings of anger.
-Consequently, in the generated image, the burning lawn is substituted with a serene scene featuring a cluster of flowers, thereby transforming the overall emotional response to awe.
-Studies Yang et al.
-(2023, 2018) have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement.
-As also seen in Figure 2, introducing fearful elements like fire or a woman in white into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments.
-However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows.
-In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses.
-Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli Baumeister et al.
-(2001); Ito et al.
-(1998), simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
+In commercialization, it enhances brand impact and captures consumer attention.
+In augmented or virtual reality, it supports therapeutic interventions for conditions such as autism and schizophrenia.
+Despite recent advances in image generation [5, 27, 37], emotion conditioning remains largely underexplored.
+We introduce the novel task of emotion-evoked image generation: given a source image and a target emotion, the goal is to generate a new image that evokes the specified emotion while preserving the semantics and structure of the original scene.
+This task is challenging, requiring both recognition of subtle emotional cues in the source image and the ability to modify them meaningfully.
+For example, as shown in Figure 1, clusters of fire flames on the water intensify the emotion of anger.
+In the generated image, they are replaced with lamps designed in cute shapes, transforming the overall emotional response to amusement.
+Studies have shown that both global and local factors influence emotional responses in images, such as overall color tones, local facial expressions, and emotion-related objects like graveyards (sadness) or balloons (amusement) [33, 32].
+As shown in Figure 2, inserting fearful elements, such as fire, into a peaceful scene can evoke negative emotions, as they trigger instinctive responses to threats.
+However, some elements, such as the sky and plants, remain emotionally neutral.
+Given that humans are generally more sensitive to negative than positive stimuli [2, 10], simply adjusting global attributes like brightness may not be enough to alter negative emotions.
+Traditional methods such as color transfer and style transfer [20, 6, 30] attempt to modify emotional tone through global appearance changes.
+However, they often fail to elicit desired emotions, as they cannot identify and edit the specific local regions responsible for emotional impact.
+In Figure 2, while darkening the scene may enhance fear (left panel), brightening a scene of a burning hillside does not effectively convey awe (right panel).
+In contrast, transforming anger into awe requires editing key emotional cues, for example, removing the fire and adding vibrant clouds to create a tranquil atmosphere.
+Emotion-evoked image generation thus requires not just global color or style adjustments, but also semantic understanding and purposeful local modification.
 Drawing inspiration from these emotion-associated factors, we propose our model, EmoEditor, for emotion-evoked image generation.
-We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties.
-(1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome.
+We develop EmoEditor based on state-of-the-art stable diffusion models and introduce three technical novelties.
+(1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process that integrates emotion-conditioned global context and local emotional cues from the source images to ensure coherence with the original content while reflecting the desired emotional outcome.
 (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions.
-(3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
-Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer Pitié et al.
-(2007); Gatys et al.
-(2015); Weng et al.
-(2023), were commonly employed to modify the emotional tone of images.
-These methods typically require reference emotional images or textual cues as additional inputs.
-However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses.
-In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
+(3) During inference, iterative emotion discrimination mechanisms are employed to select emotionally coherent images autonomously.
 Figure 2: Emotions are influenced by both global and local cues from the source images.
-We present two image generation examples showcasing the transition from positive to negative emotions in Column 1 (shaded in blue) and two examples from negative to positive emotions in Column 2 (shaded in orange).
+We present examples of transitions from positive to negative emotions (blue shading) and from negative to positive emotions (orange shading).
 In each example, the target emotions are in red, and source images are framed in green.
-Generated results from three methods are presented.
-Zoom in for better views.
-With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged Kawar et al.
-(2023); Brooks et al.
-(2023); Ruiz et al.
-(2023).
+With the rapid advancements in diffusion models, various effective image editing techniques have emerged [11, 3, 23].
 However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images.
-To address this gap, we propose a competitive baseline (Large Model Series), concatenating multiple large language and vision models.
+To address this gap, we propose a competitive baseline called Large Model Series (LMS), concatenating multiple large language and vision models.
 This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models.
-Upon comparison with existing methods through a series of human psychophysics experiments and four new quantitative evaluations, our end-to-end trained EmoEditor achieves outstanding image generation results.
-It effectively preserves the structural coherence and semantic consistency wtih the source images while eliciting the target emotion of human viewers.
-To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions.
-Unlike existing datasets on visual emotion analysis Mikels et al.
-(2005); Machajdik, Hanbury (2010); You et al.
-(2016); Peng et al.
-(2015); Yang et al.
-(2023), which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
-The main contributions of this paper are highlighted below: (1) We introduce a new and important problem of emotion-evoked image generation.
-To benchmark all methods, we establish a comprehensive framework for evaluating model performances, incorporating human psychophysics experiments and four newly introduced metrics.
-(2) We introduce EmoEditor, an emotion-evoked diffusion model, which employs a novel two-branch architecture to integrate global context with local semantic regions of source images that evoke emotional responses.
+Upon comparison with existing methods through a series of human psychophysics experiments and a new quantitative evaluation, our end-to-end trained EmoEditor achieves outstanding image generation results.
+It effectively preserves the structural coherence and semantic consistency with the source images while eliciting the target emotion of human viewers.
+To obtain training data for this task, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions.
+Unlike existing datasets on visual emotion analysis [17, 19, 33], which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
+The main contributions of this paper are highlighted below: (1) We study the problem of emotion-evoked image generation, which focuses on modifying visual content to trigger intended emotional responses while preserving scene structures and contextual coherence.
+To benchmark all methods, we establish a comprehensive framework for evaluating model performance, incorporating human psychophysics experiments and introducing a new metric.
+(2) We propose EmoEditor, an emotion-evoked diffusion model, which employs a novel two-branch architecture to integrate global context with local semantic regions of source images that evoke emotional responses.
 (3) EmoEditor learns emotional cues during end-to-end training, eliminating the need for additional emotion reference images.
@@
 Its performance is superior among all the competitive methods.
-(4) We curate the first large-scale EmoPair dataset, containing 340,000 image pairs with emotion annotations.
+(4) We curate the large-scale EmoPair dataset, containing 340,000 image pairs with emotion annotations.
 2 Related Work
 2.1 Visual Emotion Analysis
-Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses from visual data.
-In psychology, Categorical Emotion States (CES) form the basis for emotion classification models like the 8-category Mikels model Mikels et al.
-(2005).
-Early research used manually crafted features such as color and content Machajdik, Hanbury (2010), while recent advancements employ deep learning to capture emotional cues more accurately You et al.
-(2015); Yang et al.
-(2018).
-Despite progress, generating images to evoke specific emotions remains underexplored.
-We propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation to address this.
-Unlike the Affective Image Filter Weng et al.
-(2023), which requires explicit textual instructions, EmoEditor does not need text guidance to achieve desired emotions.
-Existing VEA datasets Peng et al.
-(2015); You et al.
-(2016) focus on emotion classification, including EmoSet Yang et al.
-(2023), which has 120k images with emotion labels and attributes.
+Visual Emotion Analysis (VEA) seeks to understand and predict human emotional responses from visual data.
+In psychology, Categorical Emotion States (CES) underpin classification models such as the 8-category Mikels model [17].
+Early work relied on handcrafted features like color [15], while recent methods leverage deep learning for more accurate emotion prediction [35, 32].
+However, generating images to evoke specific emotions remains underexplored.
+To address this, we introduce EmoEditor, a two-branch diffusion model that jointly learns emotion recognition and image generation.
+Existing VEA datasets [19, 36] focus on emotion classification, including EmoSet [33], which has 120k images with emotion labels and attributes.
 However, they lack image pairs with annotated emotions and text editing instructions essential for emotion-evoked image generation.
@@
 2.2 Image Editing
-Color and Style Transfer.
-Methods like PDF-TransferPitié et al.
-(2007), Neural Style TransferGatys et al.
-(2015), and CLIP-StylerKwon, Ye (2022) manipulate images using reference images, but they struggle with local region adjustments, limiting their versatility in generating emotion-evoking images.
-Our EmoEditor balances local and global manipulations without needing additional references.
-Text-to-Image Editing.
-Text-to-image editing is increasingly popular.
-Earlier studies used GANs Goodfellow et al.
-(2014) and CLIP models Radford et al.
-(2021) for image manipulation based on detailed text descriptions Patashnik et al.
-(2021); Abdal et al.
-(2022).
-Recent diffusion models like Imagic Kawar et al.
-(2023), Prompt-to-prompt Hertz et al.
-(2022), InstructPix2Pix Brooks et al.
-(2023), and DreamBooth Ruiz et al.
-(2023) generate high-quality images from text prompts but often ignore emotional aspects and rely heavily on detailed instructions.
-In contrast, our EmoEditor needs only target emotion and source image inputs, using end-to-end training to understand and edit emotional cues while preserving image structure and coherence.
+Diffusion-based Image Editing.
+Earlier studies used GANs [7] and CLIP models [21] for image manipulation based on detailed text descriptions [1].
+Diffusion models have advanced, generating high-quality images from various prompts [8, 23].
+Methods like SDEdit [16] use stochastic differential equations for editing.
+Ip2p [3] allows image editing based on human instructions.
+PnP [25] achieves fine-grained control of generated structures by manipulating spatial features and self-attention.
+CtrlNet [38] adds conditional control for better precision.
+Despite these advances, current diffusion models often overlook emotional cues needed for effective emotion editing.
+Emotion Editing.
+Color and style transfer methods like PDF-Transfer [20], Neural Style Transfer (NST) [6], and CLIP-Styler (Csty) [13] manipulate images using reference images but struggle with local adjustments, limiting their versatility in generating emotion-evoking images.
+AIF [30] reflects emotions from text to images by incorporating affective content into the image generation process.
+It relies heavily on detailed textual prompts and cannot edit image regions locally.
+EmoEdit [34] maps input images and target emotions into an embedding guided by a set of fixed learned queries, restricting emotional cues and diversity in generation to the representational capacity of these queries.
+In contrast, our EmoEditor uses end-to-end training to implicitly understand and edit emotional cues locally and globally without relying on external references.
 3 Method
 We formulate emotion-evoked image generation as a supervised learning task.
-First, we curate the EmoPair dataset containing image pairs with emotion labels and corresponding text instructions for editing (Sec.3.1).
-Next, EmoEditor is trained on this dataset (Sec.3.2).
-During inference, it takes source images and target emotions, producing generated images without the need for human-annotated text instructions or emotion reference images.
+First, we curate the EmoPair dataset with emotion-labeled image pairs and editing instructions.
+Next, EmoEditor is trained on this dataset.
+During inference, it takes source images and target emotions, producing generated images without the need for text instructions or emotion reference images.
+3.1 Curating Our EmoPair Dataset
+Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair.
+It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2p [3] annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions.
+All images from EmoPair are of size 224$\times$224.
+Figure 3 illustrates the pipeline for curating our EmoPair dataset.
+See Sec S1 for more details.
+Emotion Predictor $\mathcal{P}$.
+To construct our dataset, we use the emotion predictor $\mathcal{P}$ proposed in [33], which was trained on the largest VEA dataset, EmoSet.
+All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions.
 Figure 3: Pipeline for Curating Our EmoPair Dataset.
@@
 The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided.
-See more examples in Supp.Sec.S1.
-See Sec.3.1 for EmoPair dataset.
-3.1 Curating Our EmoPair Dataset
-Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair.
-It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2pBrooks et al.
-(2023) annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions.
-All images from EmoPair are of size 224$\times$224.
-Figure 3 shows the pipeline for curating our EmoPair dataset.
-Emotion Predictor $\mathcal{P}$.
-To better construct our dataset, we first train an emotion predictor $\mathcal{P}$ on the existing VEA dataset EmoSet Yang et al.
-(2023).
-All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions.
-$\mathcal{P}$ employs ResNet18 He et al.
-(2016) as a backbone and achieves a top-1 accuracy of 73% on EmoSet’s test set.
+Note that these instructions are used during model training to guide the learning of rich neural-symbolic embeddings, but are not used during model inference.
 EmoPair-Annotated Subset (EPAS).
@@
 EmoPair-Generated Subset (EPGS).
-To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet Yang et al.
-(2023).
-However, EmoSet lacks paired images with consistent scenes and target emotions, and it doesn’t include text editing instructions.
+To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet [33].
+However, EmoSet lacks paired images with consistent scenes and target emotions, and it does not include text editing instructions.
 Single-word prompts like “[Target Emotion]" prove ineffective in existing text-to-image models due to their limited emotional reasoning.
 Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images.
-For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape.
-Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 Brown et al.
-(2020).
+For instance, enhancing an image for excitement can vary significantly based on context, such as having a happy smile on a person or adding fireworks to the sky for a landscape.
+Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 [4].
 Human annotators rank these instructions, keeping the top ten per emotion.
-The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures (see Supp.Sec.S1).
+The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures.
+To ensure the quality of the generated images, we have human annotators providing corrections for some results (see Sec S1).
 3.2 Our Diffusion Model - EmoEditor
 Figure 4: Architecture of Our Proposed EmoEditor.
-EmoEditor is an image diffusion model, consisting of local (shaded in green) and global (shaded in blue) branches.
-The pre-trained VAE’s encoder $\mathcal{E}$ and decoder $\mathcal{D}$ remain fixed throughout training and inference.
-Exclusively employed during inference, the fixed emotion predictor $\mathcal{P}$ predicts the emotions on generated images for the iterative emotion inference.
-See Sec.3 for details.
+EmoEditor is an image diffusion model with local (shaded in green) and global (shaded in blue) branches.
+The pre-trained VAE’s encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are fixed during training and inference.
+Exclusively employed during inference, the fixed emotion predictor $\mathcal{P}$ predicts emotions on generated images for iterative emotion inference.
 Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion.
-It employs the latent diffusion model (LDM) Rombach et al.
-(2022) with three key novelties.
+It employs the latent diffusion model (LDM) [22] with three key novelties.
+Integration of Global and Local Emotion Cues.
+In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0.
+This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation.
+We use the pre-trained emotion predictor $\mathcal{P}$ to predict the source image’s emotional state $e_{s}$, representing the probability distribution of emotions for each category.
 EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios.
@@
 Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
-Integration of Global and Local Emotion Cues.
-In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0.
-This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation.
-Conversely, text inputs demand extra transformation steps, potentially introducing noise or errors.
-We introduce an emotion encoder $\tau_{\theta}$ to encode $e_{oh}$ into $e$ via fully connected layers: $e=\tau_{\theta}(e_{oh})$ (see Supp.Sec.S2.1 for details), trained from scratch to learn neuro-symbolic embeddings.
+Next, we compute the emotion editing direction $e_{dir}$ using the target emotion code $e_{oh}$ and the source emotional state $e_{s}$: $e_{dir}=e_{oh}-e_{s}$.
+This allows the model to adaptively adjust the editing direction based on the initial state of different source images.
+To map the emotion editing direction into the latent space, we introduce an emotion encoder $\tau_{\theta}$, trained from scratch to learn neural-symbolic embeddings.
+It first encodes $e_{dir}$ into an emotion embedding $e$ through fully connected layers: $e=\tau_{\theta}(e_{dir})$ (See Sec S2.1 for details).
 Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$.
 Notably, we intentionally avoid using explicit text editing instructions as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions.
-Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
+Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same emotions.
 In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$.
-Similar to the LDM Rombach et al.
-(2022), $\mathcal{E}$ is based on a pre-trained VAEKingma, Welling (2022).
+Similar to the LDM [22], $\mathcal{E}$ is based on a pre-trained VAE [12].
 The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps ($t\in T$).
@@
 Our approach loosens the constraint of precisely predicting the exact text instructions.
-This flexibility empowers our model to explore a wider array of image editing solutions, that can augment the source emotions of human viewers to target emotions.
-Consequently, our model learns to reason in a neuro-symbolic manner akin to humans, and meanwhile, preserves its creativity and capacity for exploring novel editing solutions.
-Iterative Emotion Inference.
+This flexibility empowers our model to explore a wider array of editing solutions, that can augment the source emotions of human viewers to target emotions.
+Consequently, our model learns to reason in a neuro-symbolic manner akin to humans and preserves its creativity and capacity for exploring novel editing solutions.
+Iterative Emotion Inference (IEI).
 Editing emotions is a complex process often requiring iterative image edits.
@@
 These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$.
-The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.3-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.8.
+The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.5-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.6.
 If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images.
 In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition.
-Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic.
+We use $\mathcal{P}$ to predict the emotional state of $\hat{I}$ as the new source emotional state and calculate the new emotion editing direction to update the emotion embedding $e$.
+Based on these new inputs, EmoEditor generates a new edited image for the critic.
 This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 30.
 In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
-See Supp.Sec.S2.2 for Implementation Details.
+See Sec S2.5 for more implementation details.
 4 Experiment
-We selected 504 images randomly from the EmoSet dataset Yang et al.
-(2023), ensuring they were distinct from those in our EmoPair dataset.
-According to the emotion labels of EmoSet, there are 63 images for each source emotion category.
-Subsequently, we evaluated both competitive baselines and our EmoEditor on these source images and presented the results in the following subsections.
 Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions.
-Considering emotional valence, we divided the test set into two subsets: within-valence emotion transfer and cross-valence emotion transfer.
+Considering emotional valence, we propose two tasks: within-valence emotion editing and cross-valence emotion editing.
 This division serves two purposes.
-Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween.
+Firstly, cross-valence emotion editing is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween.
 Secondly, images of the same valence may contain subtle differences challenging for humans to discern.
 For instance, an amusement park scene might evoke a mixture of excitement and amusement.
+We randomly select test images from EmoSet [33], ensuring they differ from those in our EmoPair dataset and that each source emotion category is evenly distributed.
+We then evaluate the competitive baselines and our EmoEditor on the cross-valence emotion editing task, with a total of 2,016 test images.
+We also present the results of applying our EmoEditor to more challenging emotion editing scenarios.
+To ensure a fair comparison, the input image size is standardized to 224$\times$224 across all methods.
 Baselines.
-We compare our EmoEditor with five state-of-the-art methods: (1) Color-transferPitié et al.
-(2007) (CT); (2) Neural-Style-TransferGatys et al.
-(2015) (NST).
-Both methods require an additional reference image from the target emotion category as inputs.
-We randomly selected these reference images from the EmoSet dataset.
-We also include representative text-to-image editing models (3) CLIP-StylerKwon, Ye (2022) (Csty) and (4) Ip2p Brooks et al.
-(2023).
+We compare our EmoEditor with nine methods: (1) Color-Transfer (CT) [20]; (2) Neural-Style-Transfer (NST) [6].
+Both methods require an additional reference image from the target emotion category as input.
+We randomly select these reference images from the EmoSet dataset.
+We also include representative text-driven image editing models (3) Csty [13]; (4) CtrlNet [38] (5) SDEdit [16]; (6) Ip2p [3] and (7) AIF [30].
 We provide these models with the target emotion category as a one-word text prompt, such as “excitement".
-Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing.
-To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS).
-This includes BLIP Li et al.
-(2022) for image captioning, followed by GPT-3 Brown et al.
-(2020) for text instruction generation, and Ip2p for image editing based on the instructions (See Supp.Sec.S2.3).
-[Uncaptioned image] Figure 5: Human Psychophysics Experiment Results.
-The average proportion of images that human participants prefer our EmoEditor over other methods is 56%.
-Chance is 50% (red dotted line).
-Error bars are standard errors.
-Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 33.29 7.36 NST 34.42 92.01 34.45 18.57 Csty 11.51 85.52 41.47 36.64 Ip2p 2.53 67.76 9.39 12.71 LMS 11.51 77.38 26.13 19.74 w/o $\mathcal{P}$ 5.06 69.15 19.45 14.93 w/o $L_{emb}$ 43.35 91.62 22.53 16.00 Ours 50.20 92.86 23.98 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor.
-See Sec.4 for the methods and evaluation metrics.
-See Sec.5.1 for the result analysis.
+We also compare with (8) EmoEdit [34], which performs emotion-driven image editing by retrieving semantically similar examples from an emotion dictionary.
+In addition, we concatenate existing large language and vision models and introduce the baseline (9) “Large Model Series" (LMS).
+This includes GPT-4o [9] for image captioning, followed by GPT-o4 [18] for generating reasoning-based text instructions, and Ip2p for image editing based on the instructions (See Sec S2.2).
 Human Psychophysics Experiment.
-We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)Turk (2012).
-We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials.
+We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk) [26].
+We recruit 136 participants, with each participant undergoing 180 trials, yielding a total of 24,480 trials.
 All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution.
@@
 They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them.
-The two image stimuli consists of a result generated by our EmoEditor and a randomly sampled result from the five baselines introduced above.
+The two image stimuli consist of a result generated by our EmoEditor and a randomly sampled result from the baselines introduced above.
 The trial presentation order and the binary choices for each trial are randomized.
@@
 We discard the data from those people who fail the control trials at least 17% of the time.
-See Supp.Sec.S2.4 for more details.
+See Sec S2.3 for more details.
 Evaluation Metrics.
-While CLIP Radford et al.
-(2021) is commonly used to assess text-image consistency, it doesn’t fully capture the complexity of emotions Widhoelzl, Takmaz (2024).
-Relying on CLIP for emotional alignment may overlook important nuances in images.
-Thus, we introduce four novel evaluation metrics to quantitatively assess emotion-evoked image generation performance.
-Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions.
-EMR: Similar to Goetschalckx et al.
-(2019), we also use our emotion predictor $\mathcal{P}$ to assess the emotional category of generated images and determine their alignment with target emotions.
-We acknowledge that using $\mathcal{P}$ as an evaluator may introduce favorable biases toward our model.
-However, the absence of existing emotion evaluators leaves us no choice but to rely on $\mathcal{P}$.
-ESR: $\mathcal{P}$ generates emotion probability distributions $e_{g}$ and $e_{s}$ for generated and source images.
-We compute ESR using Kullback-Leibler Divergence distance $KLD(\cdot)$, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$.
-$e_{oh}$ is the binary one-hot target emotion vector, indicating the impact of edits on generated images transitioning from the source emotion to the target emotion.
-Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) assess the structural coherence and semantic consistency between source and generated images.
-ENRD: Using Grad-CAM Selvaraju et al.
-(2017) with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify emotionally neutral regions on source images, valued at 0.
-Then, we compute the pixel-level L1 distance between these regions on source and generated images.
-ESS: Employing the Canny edge detection algorithm with thresholds of 200 and 500 on both source and generated images, we determine the L1 norm between their edge maps to quantify structural disparities.
+While CLIP [21] assesses text-image consistency, it does not fully capture emotional complexity [31].
+Thus, we introduce Emotional Structure Matching Index (ESMI): $ESMI=\alpha\times S_{str}+(1-\alpha)\times S_{emo}$ where $\alpha$ is a weighting coefficient and set to 0.5.
+ESMI evaluates emotion-evoked image generation from two aspects: evoking specific emotions and maintaining structural and semantic consistency between the source and generated images.
+Evoking Targeted Emotions.
+To assess how well the generated image expresses the target emotion, we use $\mathcal{P}$ to extract emotional distributions $e_{s}$ and $e_{g}$ from the source and generated images respectively.
+Next, we calculate the Kullback-Leibler Divergence between these distributions and the target emotion’s one-hot encoding $e_{oh}$, we get $D_{src}=KLD(e_{s},e_{oh})$ and $D_{gen}=KLD(e_{g},e_{oh})$.
+Ideally, the generated image should emotionally be closer to the target, that is, $D_{gen}<D_{src}$.
+To avoid the influence of outliers or extreme values, we define the raw emotional improvement as $\Delta D=\max(0,D_{src}-D_{gen})$.
+To further normalize this improvement relative to the original emotional distance, we compute the emotional evocation score: $S_{emo}=\Delta D/(D_{src}+\epsilon)$ where $\epsilon$ is a small constant to prevent division by zero.
+This score ranges from 0 to 1, with higher values indicating a greater emotional shift toward the target emotion relative to the source.
+Preserving Original Structures.
+We define two regions: the Emotion-Evoking Region ($R_{emo}$), which is responsible for triggering the source emotion (e.g.
+fire for anger) and is expected to be altered, and the Emotion-Neutral Region ($R_{neu}$), which should remain unchanged.
+We measure structural consistency by computing the pixel-level L1 differences between the source and generated images in these regions, denoted as $L_{emo}$ and $L_{neu}$.
+The structural preservation score is calculated as: $S_{str}=L_{emo}/(L_{emo}+L_{neu}+\epsilon)$ where $\epsilon$ is a small constant to avoid division by zero.
+This score reflects the proportion of changes in $R_{emo}$ relative to all changed pixels, with higher values indicating better preservation of neutral areas.
+We adopt two strategies to define $R_{emo}$: (1) CAM-based: $R_{emo}$ is obtained by applying Grad-CAM [24] to $\mathcal{P}$ and binarizing the resulting heatmap.
+(2) HA-based: $R_{emo}$ is derived by averaging manually annotated emotional regions of human annotators.
+See Sec S2.4 for more details.
+ESMI(%) ↑ CT NST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
+CAM-based 28.33 47.48 34.33 32.94 35.91 27.37 29.71 32.94 35.40 51.56
+HA-based 26.68 46.00 32.77 31.53 34.27 25.72 28.31 37.53 33.76 49.99
+Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods.
+The CAM-based metric computes ESMI using $R_{\text{emo}}$, derived by binarizing the Grad-CAM output from the Emo Predictor.
+The HA-based metric computes ESMI using $R_{\text{emo}}$, defined as the average of human-annotated regions.
+Best is in bold.
+Larger (↑) is better.
+Figure 5: Results from Human Psychophysics Experiments.
+The proportions of images that human participants prefer our EmoEditor over all other methods on the x-axis are presented.
+Chance is 50% (red dotted line).
+Error bars are standard errors.
+Variant Input of $\tau_{\theta}$ IEI $L_{emb}$ ESMI(%)↑
+text $e_{oh}$ $e_{dir}$
+1 5.85
+2 50.24
+3 7.36
+4 8.51
+5 47.36
+6 49.80
+7 (ours) 51.56
+Table 2: Quantitative Evaluation in CAM-based ESMI for Model Variants of our EmoEditor.
+A tick indicates the inclusion of a component in the corresponding model variant, while a blank cell indicates its absence.
+The last row represents our full EmoEditor model.
+Best results are in bold.
+Larger (↑) is better.
 5 Results
-5.1 Quantitative Evaluation in Cross-Valence Scenarios
-Figure 5 shows human psychophysics experimental results (see Supp.Sec.S3.1 for more results).
-If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks.
-Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec.4.
+5.1 Evaluation in Cross-Valence Scenarios
+Quantitative Evaluation with SOTA.
+Figure 5 shows human psychophysics experimental results (see Sec S3.1 for more results).
+If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with a 50% chance in the “two-alternative forced choice" tasks.
 Observing Figure 5, we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines, indicating its proficiency in evoking target emotions.
-This is supported by EmoEditor’s highest EMR and ESR scores in Table 1.
-Additionally, EmoEditor achieves moderately lower ENRD and ESS, suggesting maintained structural and semantic coherence in the generated images.
-Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Supp.Fig.S10 for more examples.
-From Figure 5, we noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than global color and brightness adjustments.
-CT also has lower EMR and ESR than our EmoEditor in Table 1.
-As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores.
-Similarly, our EmoEditor is significantly favored over Ip2p by human participants, indicating that current text-to-image diffusion models face challenges in understanding emotional cues and generating effective text instructions for editing emotion-evoked images.
-Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions.
-Additionally, our EmoEditor also outperforms Large Model Series (LMS) based on human preference scores in Figure 5, 39% in EMR, 15% in ESR, 2 in ENRD, and 3 in ESS in Table 1.
-This suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
-Neural-Style-Transfer (NST) and CLIP-Styler (Csty) excel in inducing negative emotions through distortion and irregular textures on source images.
-They are preferred in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.Fig.S8).
-However, these methods render images challenging to comprehend, indicated by higher ENRD and ESS scores in Table 1, suggesting the distorted images lose structural similarities and semantic content.
-In contrast, our EmoEditor outperforms both methods in generating positively valenced images, with higher preference scores in the lower quadrants of the confusion matrix (see Supp.Fig.S8).
-In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance.
-This leads to lower EMR and ESR scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions, and executing iterative refinements during inference.
-We also perform ablation studies on different max iterations for iterative emotion inference, showing that this iterative process improves the model’s performance, achieving stable and high-quality emotionally evocative images (see Supp.Sec.S3.3).
-5.2 Visualization of Emotional Image Generation across Valence
-Figure 1 and 6 present visualizations of generated images for all competitive methods in cross-valence scenarios (see Supp.Sec.S3.2 for more results).
-Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation.
-Neural Style Transfer heavily relies on randomly selected reference images.
-While it produces artistic textures aligned with target emotions, it lacks in preserving essential semantic content in the source images.
-CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Column 4.
-Ip2p demonstrates limited image generation abilities with single-word emotion prompts.
-For example, the fire in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 1, Column 5).
-Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions.
-This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 2, Column 6.
-In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible.
-For instance, in Row 1, Column 7, our method does not change the subject of the fire but turns the fire in the woods into a bonfire and generates a beautiful night sky in the background, thus evoking contentment.
-In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
+Complimentary to psychophysics experiments, Table 1 presents quantitative results in CAM-based and HA-based ESMI, with our EmoEditor achieving the highest scores in both metrics.
+Notably, for each method, the scores remain consistent across the two evaluation metrics, indicating that the emotion predictor $\mathcal{P}$ identifies emotion-evoking regions in a manner aligned with human annotations.
+This consistency further supports the reliability and validity of the CAM-based ESMI metric.
+Our EmoEditor significantly outperforms CT, demonstrating that emotion-evoked image generation requires more than global color adjustments.
+While NST and Csty can induce negative emotions through distortions and irregular textures, they often reduce image interpretability (see Fig S10).
+CtrlNet and SDEdit produce abstract textures that may evoke emotion but often fail to preserve the source image’s semantics.
+Human evaluations also prefer EmoEditor over Ip2p, underscoring the limitations of current diffusion models trained for general image generation when applied to emotion-driven tasks.
+Although both AIF and EmoEdit are designed for emotional editing, our EmoEditor achieves higher human preference and ESMI scores.
+AIF applies global filters, resulting in less targeted changes.
+EmoEdit maps image-emotion pairs into an embedding, but the limited and biased emotion dictionary leads to repetitive edits and poor performance on strongly emotional inputs.
+Finally, our EmoEditor outperforms LMS in both human preference and ESMI, suggesting that simply chaining large models may amplify biases and errors.
+Emotion-evoked generation demands a nuanced understanding of both the source image and target emotion, making end-to-end learning crucial.
+Figure 6: Visualization of Generated Images from Different Methods.
+Target emotion is highlighted in red and source image is framed in green.
+The blue number below each image is its CAM-based ESMI score, with higher values being better.
+More examples in Sec S3.2.
 Figure 7: Our EmoEditor can generalize to more challenging emotion editing scenarios.
-(a) Three examples are showcased, highlighting our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images.
-(b) Two examples demonstrate how our EmoEditor can transform neutral real-world images to evoke either positive or negative emotions.
+(a) Same-Valence Editing highlights our EmoEditor’s ability to produce images evoking emotions of the same positive valence as the source images.
+(b) Neutral-Valence Editing demonstrates how our EmoEditor can transform neutral real-world images to evoke positive or negative emotions.
+(c) Multi-Emotion Editing demonstrates its capability to generate images that evoke a wider range of emotions, providing more nuanced and diverse visual outputs.
 Source images are framed in green.
 Target emotions are in red.
-See Sec.5.3 for the result analysis.
+Model Analysis.
+We conduct the model analysis and present the results in Table 2.
+In Row 1, the model variant takes only text as input and follows the structure of Ip2p.
+Although it is fine-tuned on our EmoPair dataset with the pre-trained CLIP model as $\tau_{\theta}$, the performance is the lowest among all variants, indicating that text guidance is insufficient.
+Adding IEI (Row 2) results in notable improvements over the version without IEI (Row 1), confirming its contribution to enhancing emotional consistency.
+Next, replacing the text input with a one-hot encoding of the target emotion $e_{oh}$ (Row 3) leads to a better result, highlighting the advantage of structured emotion encoding over text.
+Further improvement is observed when replacing $e_{oh}$ with an emotion editing direction $e_{dir}$ (Row 4), suggesting that incorporating information about the source emotion enables more precise and effective editing.
+Adding the alignment loss $L_{emb}$ (Row 6) improves performance over the variant without it (Row 5), confirming its complementary role in promoting emotionally consistent generation.
+Our final model (Row 7) integrates $e_{dir}$, IEI, and $L_{emb}$, enabling emotion-aware editing based on the source-target gap.
+This configuration achieves the highest score, demonstrating its superior emotional precision and generative diversity.
+Visualization of Emotional Image Generation across Valence.
+Figure 1 and 6 present visualizations of generated images in cross-valence scenarios.
+CT primarily replicates tonal characteristics from a reference image, often failing to significantly enhance emotions.
+NST relies on randomly selected references, producing artistic textures aligned with target emotions but losing essential semantic content.
+AIF, dependent on emotional text descriptions, applies a global filter effect but struggles with localized editing.
+EmoEdit relies on fixed learned queries, leading to limited, context-incongruent edits (e.g., always adding balloons for amusement; Figure 1) and failing to edit local emotional cues such as fire flames or fireworks (Figure 6).
+Even with a richer vocabulary, such as LMS, which integrates three large models, it still struggles to generate desirable outputs.
+In contrast, our EmoEditor takes only a source image and target emotion as input, yet generates creative and emotionally aligned results while preserving structural and semantic coherence.
+By learning an emotion editing direction, it understands the source-target gap and applies precise edits.
+In Figure 6, where fire and fireworks dominate the original emotion, EmoEditor adapts its edits to the target: extinguishing flames and adding mountains to evoke excitement (a), or removing fireworks and darkening tones to convey sadness (b).
+In Figure 1, it replaces fire flames with lamps to elicit amusement.
+Notably, even when handling similar anger-inducing elements like fire, our EmoEditor produces diverse, context-aware edits tailored to the source image without any manual design, demonstrating a nuanced understanding of emotional cues.
+We also show CAM-based ESMI scores (in blue) for each result.
+Our EmoEditor achieves the highest scores, demonstrating its ability to accurately identify and effectively modify emotion-evoking regions in source images, striking a balance between evoking the target emotion and preserving the original structure.
 5.2 Generalization to Real-world Scenarios
-Editing images within the same emotional valence presents greater challenges compared to cross-valence emotion image editing due to subtle differences in emotional cues.
-Figure 7(a) showcases our EmoEditor’s results within the same valence.
-Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes.
-Example 1 transitions from awe to contentment, depicting a sunset scene with afterglow.
-Example 2 generates a blooming white flower with additional purple flowers to evoke excitement.
-Example 3 transforms a lively spring scene into a tranquil night scene with moonlight to evoke awe.
+Editing images with the same emotional valence is more challenging due to subtle emotional cues.
+Figure 7a demonstrates EmoEditor’s ability to handle such tasks, as it removes a parachutist to reveal a breathtaking landscape, shifting excitement to awe.
 In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all.
-We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset Lin et al.
-(2014), commonly used for object detection.
-These images depict everyday life scenes with neutral emotions.
-Our EmoEditor applies positive and negative emotion transformations to these source images.
-Figure 7(b) illustrates the generated results.
-In Example 1, to inject amusement into an office space, our EmoEditor embellishes the walls with delightful watercolor paintings and infuses lively and vibrant colors into the room furniture.
-Conversely, to evoke fear, our model creates an aged desk and dimly lit walls.
-In Example 2, for an image featuring a courtyard with a teddy bear observing the garden, our method enriches the scene by adding a beautiful flower bed, creating an awe-inspiring effect.
-Conversely, to evoke feelings of disgust, our EmoEditor removes the flowers from the courtyard and transforms the view into a dilapidated scene with shabby clothes hanging on the walls and trash bags in the corner.
-It is crucial to emphasize that our EmoEditor is not trained in these neutral scenarios.
-All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
+We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset [14].
+These images depict scenes of everyday life with neutral emotions.
+EmoEditor applies positive and negative emotion transformations to these source images.
+It enhances the lushness of a forest to evoke contentment and transforms it into a barren landscape for disgust (Figure 7b).
+Notably, EmoEditor generates these results purely from its creativity, without any human-annotated editing instructions or scene interpretations.
+Furthermore, EmoEditor can manage multi-emotion editing with the use of emotion encoding as input.
+We extend the single-target emotion scenario by randomly adding one additional emotion of the same valence as the target and then use EmoEditor for image generation.
+In Figure 7c, EmoEditor generates a majestic mountain in the flame area of the source image to evoke awe.
+After adding contentment, it enriches the area with dense forests.
 6 Discussion
-We present a new and important problem of emotion-evoked image generation.
-To address this, we present EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics.
+We introduce the challenging problem of emotion-evoked image generation and propose EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics.
 Moreover, we contribute EmoPair dataset for model training.
-To benchmark all the methods, we introduce new evaluation metrics and establish standard protocols in emotion understanding, visualization, and reasoning.
-While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images.
-See Supp.Sec.S3.4 for details.
-The problem of emotion-evoked image generation introduced here opens doors for research across AI, psychology, design, arts, and neuroscience.
-However, EmoEditor, like any technology, carries inherent risks.
-Misusing it, such as manipulating images to evoke undesirable emotions, could harm individuals’ mental well-being or mislead public feelings and opinions.
-Supplementary Material
+To benchmark all methods, we introduce a new evaluation metric that effectively measures the balance between evoking target emotions and maintaining the structural and semantic integrity of the source image.
+While EmoEditor excels in quantitative performance and visual results, it struggles with fine details on faces (See Sec S3.4).
+While we adopt the 8-category Mikels model, real-world emotions are far more nuanced and complex.
+In particular, differentiating between positive emotions tends to be more challenging than distinguishing negative ones.
+Additionally, EmoEditor carries risks of misuse, such as evoking undesirable emotions, which could harm mental well-being or mislead public sentiment.
+All of these aspects warrant further investigation in future research.
 Appendix S1 EmoPair Dataset
-In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2pBrooks et al.
-(2023) and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
-For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 Brown et al.
-(2020).
+In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2p [3] and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
+For the EmoPair-Annotated Subset (EPAS), we use $\mathcal{P}$ to categorize the source and target images from the Ip2p dataset into eight different emotions.
+We remove samples where the emotional labels of the source and target images are consistent, then use the emotion labels to augment the Ip2p dataset, generating 331,595 image pairs.
+Figure S1: Instruction examples of EmoPair-Generated Subset (EPGS).
+For each emotion category, we retained the top ten text instructions based on the rankings determined by human annotators according to the efficacy of each emotion category.
+We utilized these instructions for Ip2p in image editing to generate target images capable of evoking the desired emotions, thereby constructing our EPGS.
+For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 [4].
 Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten.
@@
 We employ the following selection criteria to control the quality of generated image pairs of EPGS: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions.
-(2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) Wang et al.
-(2004) and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al.
-(2018) for filtering the remaining outcomes.
+(2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) [29] and Learned Perceptual Image Patch Similarity (LPIPS) [39] for filtering the remaining outcomes.
 SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences.
-Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1.
+Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: $0.60.3<SSIM(x,\hat{x})<$ 0.6 and $LPIPS(x,\hat{x})>0.1$.
 Ultimately, EPGS retains 6,949 image pairs.
-To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
-Figure S1: Instruction examples of EmoPair-Generated Subset (EPGS).
-For each emotion category, we retained the top ten text instructions based on the rankings determined by human annotators according to the efficacy of each emotion category.
-We utilized these instructions for Ip2p in image editing to generate target images capable of evoking the desired emotions, thereby constructing our EPGS.
+To ensure the quality of EPGS, we have human annotators re-annotate 300 images.
+The annotators view both the source and edited images, and we ask them to select the one that better evokes the target emotion.
+Since we need the target image to evoke the desired emotion more effectively than the source image, if the annotators choose the source image, we swap the source and target images in the original dataset to create a new subset.
+Each image is reviewed by three annotators, and the final subset is determined by majority vote.
+We then use this subset to fine-tune our EmoEditor.
+To provide a comprehensive overview of our dataset, additional image pair examples are presented in Figure S3 and S3.
 Figure S2: Sample of EmoPair-Annotated Subset (EPAS).
@@
 S2.1 Emotion Encoder
-Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming one-hot encoded emotion input vectors into the structured emotion embedding $e$.
+Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$.
+We represent the target emotion as a one-hot encoding $e_{oh}$ and use the emotion predictor $\mathcal{P}$ to calculate the emotional state of the source image $e_{s}$.
+Then, we compute the emotion editing direction $e_{dir}$ by calculating the difference between $e_{oh}$ and $e_{s}$.
+Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming emotion editing direction $e_{dir}$ into the structured emotion embedding $e$.
 The architecture of the model is composed of a series of fully connected layers that progressively increase the dimensionality of the input.
-Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$.
 The network begins with an input layer of size 8, which is then passed through a sequence of fully connected layers with increasing sizes: 256, 512, and 768 neurons, respectively.
@@
 Figure S4: The network structure of our emotion encoder $\tau_{\theta}$.
-S2.2 Implementation Details
-Follow Wang et al.
-(2022), we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p Brooks et al.
-(2023) weights.
-Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$.
-The frozen emotion predictor $\mathcal{P}$ is only used during inference.
-We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer.
-The max time step $T$ is set to 1,000.
-S2.3 Baseline: Large Model Series
+S2.2 Baseline: Large Model Series
 Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing.
 To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS).
-This includes BLIP Li et al.
-(2022) for image captioning, followed by GPT-3 Brown et al.
-(2020) for text instruction generation, and Ip2p for image editing based on the instructions.
+This includes GPT-4o [9] for image captioning, followed by GPT-o4 [18] for generating reasoning-based text instructions, and Ip2p for image editing based on the instructions.
+Figure S5: Workflow of Large Model Series (LMS).
+We decompose emotion-evoked image generation into three steps: employing GPT-4o for image understanding, utilizing GPT-o4 for instruction generation, and employing Ip2p for image editing.
+We refer to this approach of chaining multiple large models as the Large Model Series.
 Figure S5 shows the workflow of LMS.
-First, we employ BLIP to comprehend the source image, generating an image caption corresponding to the source image.
+First, we employ GPT-4o to comprehend the source image, generating an image caption corresponding to the source image.
 Subsequently, based on the generated caption, we utilize the sentence structure “There is an image with [Image caption].
 To make me feel [Target Emotion], how to change the image?
-Give one instruction that starts with an action." to query GPT-3.
-Following this, GPT-3 generates an instruction to guide Ip2p in editing the source image, resulting in the final output.
-Figure S5: Workflow of Large Model Series (LMS).
-We decompose emotion-evoked image generation into three steps: employing BLIP for image understanding, utilizing GPT-3 for instruction generation, and employing Ip2p for image editing.
-We refer to this approach of chaining multiple large models as the Large Model Series.
-S2.4 Human Psychophysics Experiment
+Give one instruction that starts with an action." to query GPT-o4.
+Following this, GPT-o4 generates an instruction to guide Ip2p in editing the source image, resulting in the final output.
+S2.3 Human Psychophysics Experiment
 Figure S6(a) shows the MTurk experiment schematic and Figure S6(b) shows the experiment instruction pages given to the participants.
@@
 Considering individual emotional response variability, participants are allowed a maximum of one incorrect choice in these dummy trials.
-Subjects exceeding an error rate of 1/6 are excluded, resulting in 11,400 trials.
-The outcomes of the 6 dummy trials are also excluded from the final statistics.
+Subjects exceeding an error rate of 1/6 are excluded, resulting in 24,480 trials.
+The outcomes of the 6 dummy trials are also excluded from the final result analysis.
 Figure S6(c) provides an example of a dummy trial.
 (b) Each participant can only take part in the experiment once.
-(c) Image pairs used in the entire experiment are drawn from the 2,016 result pairs generated by our model and five other SOTA methods.
+(c) Image pairs used in the entire experiment are drawn from the 2,016 result pairs generated by our model and other SOTA methods.
 Participants view randomly sampled pairs from these 2,016 pairs, and all trials are presented in a random order.
@@
 Participants are presented with a set of images and a target emotion.
-They must choose between two images, one generated by our model and the other by five state-of-the-art methods, selecting the one that more effectively evokes the target emotion.
+They must choose between two images, one generated by our model and the other by 9 state-of-the-art methods, selecting the one that more effectively evokes the target emotion.
 (b) Instructions of the MTurk experiment.
@@
 (d) Real trial example.
-Appendix S3 Results
-S3.1 Quantitative Evaluation in Cross-Valence Scenarios
-We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)Turk (2012) (Online).
-We recruited 417 participants, each undergoing 100 trials.
-Additionally, we conducted in-lab experiments, involving 10 volunteers.
-Similarly, each participant reviewed 100 trials.
-The consistency between online and in-lab experimental results demonstrates the superiority of our approach in producing results that evoke the target emotion compared to other state-of-the-art methods.
+S2.4 Evaluation Metrics: Structural Score
+Emotion-evoked image generation needs to be evaluated from both emotional induction and structural preservation perspectives.
+Traditional metrics, such as SSIM, are inadequate for structural preservation evaluation because they do not consider emotional cues.
+The regions in the image that evoke emotions are called Emotion-Evoking Regions ($R_{emo}$).
+These areas need to be altered to change the source emotion and induce the target emotion.
+Conversely, some regions in the image lack significant emotional cues and are referred to as Emotion-Neutral Regions ($R_{neu}$).
+These regions should be preserved as much as possible during the editing process.
+Grad-CAM [24] can visualize which parts of the image most influence the model’s decision.
+Therefore, we use Grad-CAM with the emotion predictor $\mathcal{P}$ to identify Emotion-Evoking Regions ($R_{emo}$) in the source images.
+Figure S7 shows the visualization of the Emotion-Evoking and Emotion-Neutral Regions.
+Using Grad-CAM [24] with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify Emotion-Evoking Regions (valued at 1) and Emotion-Neutral Regions (valued at 0) on source images.
+To evaluate changes, we calculate pixel-level L1 differences between the source and generated images for these regions, denoted as $L_{emo}$ and $L_{neu}$.
+The structural score is: $S_{str}=L_{emo}/(L_{emo}+L_{neu}+\epsilon)$ where $\epsilon$ is a small constant to avoid division by zero.
+This score reflects the proportion of changes in $R_{emo}$ relative to all changed pixels, with higher values indicating better preservation of neutral areas.
+Figure S7: Visualization Explanation of the Emotion-Evoking and Emotion-Neutral Regions.
+The source image is framed in green.
+Using Grad-CAM [24] with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify Emotion-Evoking Regions (valued at 1) and Emotion-Neutral Regions (valued at 0) on source images.
+Considering that Grad-CAM may be inaccurate or unstable due to the limitations of the emotion predictor $\mathcal{P}$, we invited 46 participants to annotate the emotion-evoking regions of 504 source images.
+Each worker annotated 50 trials, resulting in a total of 2,300 trials.
+Additionally, we included 6 dummy trials, which were the same as those used in our human psychophysics experiment (Figure S6(c)), to ensure annotation quality.
+Results from workers who failed more than once on the dummy trials were discarded.
+Ultimately, each image was annotated by at least three individuals.
+We then averaged all the annotations and binarized the results to generate a human-annotated weight map.
+Figure S8 shows example annotations and a comparison with Grad-CAM.
+Figure S8: Emotion-Evoking Region Annotated by Human.
+Each image has been annotated by at least three human annotators who identified the emotion-evoking regions.
+We averaged all the annotations and binarized the result to generate a human-annotated weight map.
+Row 2 presents both the weight map and the heatmap computed using Grad-CAM, showing that the two are largely consistent.
+S2.5 Implementation Details
+Following [28], we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p [3] weights.
+Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$.
+The frozen emotion predictor $\mathcal{P}$ is only used during inference.
+We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer.
+The max time step $T$ is set to 1,000.
 Figure S9: Preference of Our Method Over SOTA Methods.
-We conducted both online experiments (417 participants) and in-lab experiments (10 participants).
-The left column represents the online results, while the right column displays the in-lab results.
 The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions.
@@
 Error bars are standard errors.
-Figure S7 illustrates the preference of our method over other state-of-the-art methods in both online and in-lab experiments.
-From the overall results (Row 1), the in-lab experiment demonstrates a higher preference compared to the online experiment.
-This suggests that our method’s effectiveness in evoking the target emotion is more pronounced in a laboratory setting.
-Note that our in-lab and online experiments had identical setups.
-The observed difference may stem from the controlled environment provided by the lab conditions, where participants were more focused, reducing randomness and facilitating a more accurate assessment of emotion generation.
-This is further supported by the dummy trial success rates, with a 100% pass rate in the in-lab experiment compared to only 27% in the online experiment.
-Again, we emphasize that we only use the online data for result analysis, where the participants passed the dummy tests with less than 1/6 error rates.
-Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-TransferGatys et al.
-(2015) (NST) and CLIP-StylerKwon, Ye (2022) (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes.
-Additionally, our EmoEditor surpasses Color-transferPitié et al.
-(2007) (CT) and Ip2pBrooks et al.
-(2023) in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions.
-This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones.
-Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
-In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S8 and S9.
+Appendix S3 Results
+S3.1 Quantitative Evaluation in Cross-Valence Scenarios
+We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)[26] (Online).
+We recruit 136 participants, with each participant undergoing 180 trials, yielding a total of 24,480 trials.
+Figure S9 illustrates the preference of our method over other state-of-the-art methods, demonstrating the superiority of our EmoEditor in producing results that evoke the target emotion compared to other state-of-the-art methods.
+From the overall results (Row 1), we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines.
+Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-Transfer (NST) [6] and CLIP-Styler (Csty) [13] perform comparably to our method in generating images that evoke negative emotions, but struggle with positive outcomes.
+This indicates that negative emotions are easily influenced by factors such as chaotic textures and overall tones, whereas positive emotions are more challenging and require an understanding of the source image along with effective adjustments.
+Our EmoEditor significantly outperformed Color Transfer (CT) [20] and AIF [30] in all scenarios.
+This confirms that emotion-evoked image generation involves more than just global color and brightness adjustments.
+Additionally, our EmoEditor surpasses SDEdit [16] and Ip2p [3] in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions.
+This highlights challenges in current diffusion models for understanding and generating emotion-evoked images and indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones.
+For both positive and negative emotion generation, our method is preferred over EmoEdit in human evaluations.
+Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable.
+While LMS achieves slightly higher scores in positive emotion generation, the margin is narrow and largely influenced by its reliance on powerful external reasoning modules.
+In contrast, our method offers a compact and fully end-to-end solution that generalizes well across both positive and negative emotions, without the need for multi-stage processing or large-scale language models.
+Moreover, these language models are often proprietary and equipped with safety constraints, which limit their applicability in privacy-sensitive contexts.
+In addition, we show the confusion matrix of different methods in Figure S10.
 The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories.
-From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer Gatys et al.
-(2015) and CLIP-Styler Kwon, Ye (2022) more in the manipulation of emotions from positive to negative.
+From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer (NST) [6] and CLIP-Styler (Csty) [13] more in the manipulation of emotions from positive to negative.
 This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image.
-In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images compared to these two approaches.
-Figure S8: Confusion matrix of different methods (Online).
+In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images.
+Overall, for generating positive-valence images, our EmoEditor achieves the best results in generating images that evoke awe and contentment, but it is slightly less effective in generating images that evoke excitement.
+This may be because excitement is a more dynamic and complex emotion, requiring a more intricate combination of visual elements.
+Figure S10: Confusion matrix of different methods.
 The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs.
-The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods.
-The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods.
+The top-left matrix presents the average overall results of our EmoEditor compared to the other SOTA methods.
+The remaining nine matrices show individual comparisons between our EmoEditor and each of the nine SOTA methods.
 Average preferences for our EmoEditor and other methods are indicated above each matrix.
@@
 Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
-Figure S9: Confusion matrix of different methods (In-lab).
-The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs.
-The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods.
-The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods.
-Average preferences for our EmoEditor and other methods are indicated above each matrix.
-The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion.
-Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
+Table S1 presents the quantitative evaluation of generated images for all competitive methods.
+Our EmoEditor achieves the best performance in both $S_{emo}$ and ESMI.
+Although Ip2p scores highest on $S_{str}$, this is because it makes minimal changes to the original images in most cases, as reflected by its low $S_{emo}$ scores and visual results.
+In contrast, our method effectively edits the emotion-evoking regions in the source images, achieving a balance between evoking the target emotion and preserving the structure, which is supported by the highest ESMI score.
+Metric CT NeurST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
+$S_{emo}$(%)↑ 27.70 65.50 38.54 36.34 39.16 19.92 31.42 46.88 38.77 69.89
+$S_{str}$(%)↑ 28.96 29.47 30.11 29.55 32.65 34.82 28.00 31.60 32.03 33.22
+ESMI(%)↑ 28.33 47.48 34.33 32.94 35.91 27.37 29.71 39.24 35.40 51.56
+Table S1: Quantitative Evaluation of Generated Images for All Competitive Methods. Best in bold. Larger (↑) is better.
+Figure S11: Visualization of Generated Images from Different Methods (from negative to positive). The target emotion is highlighted in red and the source image is framed in green.
+Figure S12: Visualization of Generated Images from Different Methods (from positive to negative). The target emotion is highlighted in red and the source image is framed in green.
 S3.2 Visualization of Emotional Image Generation across Valence
-Figure S10 illustrates the visualizations of images generated by all methods across various scenarios.
-Color-Transfer Pitié et al.
-(2007) (CT) proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image.
-For instance, in Row 1, Column 2, it enhances brightness but fails to evoke amusement as it does not alter elements causing sadness, such as the crying woman.
-Neural-Style-Transfer Gatys et al.
-(2015) (NST) heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image.
-For example, in Row 4, Column 3, it generates multiple differently colored blocks but fails to evoke the emotion of sadness.
-CLIPstyler Kwon, Ye (2022) (Csty) introduces inexplicable textures.
-Ip2p Brooks et al.
-(2023), due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities, as seen in Column 5.
-Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion, as evidenced by Row 1, Column 6, where it not only fails to change the crying woman but amplifies the crying details.
-In addition, it often greatly changes the scene of the source image to achieve the goal of changing the emotion, such as Row 2, Column 6, it produces a building that looks very technological but quite different from the source image.
+Figure S11 illustrates the visualizations of images generated by all methods for the task of generating positive-valence images from negative emotions.
+Color-Transfer (CT) [20] proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image.
+For instance, in example(b), it enhances brightness but fails to evoke awe as it does not alter elements causing fear.
+Neural-Style-Transfer (NST) [6] heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image.
+For example, in example(a), it generates multiple differently colored blocks but fails to evoke the emotion of excitement.
+CLIPstyler (Csty) [13], SDEdit [16], and CtrlNet [38] often introduce inexplicable textures, losing the semantics and structure of the source images.
+Ip2p [3], due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities.
+AIF [30] heavily relies on text descriptions for understanding emotions and can only perform global edits, similar to applying a filter.
+As a result, it fails to achieve effective local edits to evoke the target emotion.
+EmoEdit [34] has difficulty accurately identifying and localizing strongly emotion-evoking regions in source images, resulting in ineffective or imprecise edits.
+In example (a), Ip2p, AIF, and EmoEdit fail to understand that the fear in the source image is caused by a strange shadow.
+As a result, they do not alter the shadow and thus cannot counteract the fear.
+Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion.
+In addition, it often greatly changes the scene of the source image to achieve the goal of changing the emotion.
+In example (c), it fails to recognize that the large flames in the background are the primary source of the original emotion, and instead generates an image where the person remains in front of the fire, merely adjusting the flame colors to be more vibrant.
+Such changes are insufficient to alter the original emotional impact.
 In contrast, our EmoEditor, requiring only the target emotion and source image inputs, can generate highly creative images that evoke the target emotion while striving to maintain scene structure and semantic coherence.
-For example, in Row 1, Column 7, it changes the woman’s pose while retaining the main scene and character from the source image, generating a cartoon-like image that aligns with amusement.
-In Row 2, Column 7, our EmoEditor generates technological results that make people feel awe while almost maintaining most of the structure of the source image.
-Figure S10: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green.
-S3.3 Ablation Study of Different Iterations
-Due to the abstract and complex nature of emotions, we propose an iterative emotion inference approach.
-The maximum number of iterations significantly impacts the quality of images generated by the model.
-To assess this effect, we vary the maximum number of iterations from 1 to 30 and evaluate the results using four metrics.
-Figure S11 presents our experimental findings.
-The Emotion Matching Ratio (EMR) shows a steady increase from 0% to approximately 50% over 30 iterations, indicating continuous improvement in the model’s ability to generate images that evoke the target emotions.
-The Emotion Similarity Ratio (ESR) rises rapidly in the initial iterations, stabilizing around 90%.
-This suggests that the model quickly becomes proficient at producing images that closely match the target emotional content, with minimal improvements after the initial iterations.
-The Emotion-Neutral Region Deviation (ENRD) also increases rapidly at first, plateauing around 22.5.
-This pattern indicates that the deviation in neutral regions of the generated images becomes consistent after a few iterations, reflecting the model’s stability in preserving non-emotional aspects.
-The Edge Structure Similarity (ESS) follows a similar trend, rising quickly to around 16 and then stabilizing.
-This stability in ESS reflects the model’s ability to maintain structural coherence and semantic consistency between the source and generated images after the initial iterations.
-In summary, these metrics demonstrate that the iterative process enhances the model’s performance significantly in the initial stages, reaching a stable state of high quality and consistency in generating emotionally evocative images.
-Figure S11: Performance Metrics Across Different Iterations.
+In example(a), our EmoEditor understands that the shadow in the source image is the key factor triggering fear.
+Instead of altering the shadow’s shape, it generates a sunset sky background, transforming the shadow into a cloud in the sky, which successfully evokes the target emotion of excitement.
+In example(b), it replaces a Halloween pumpkin with a castle to elicit the target emotion of awe.
+In example(c), it replaces the background flames with a red sunset in an attempt to evoke contentment.
+Notably, the generation process involves no manual intervention, showcasing the imagination and creativity of our model.
+Additionally, Figure S12 shows the generation results of different methods for transitioning from positive to negative emotions.
+Compared to generating positive-valence images, CT and AIF perform better in generating negative-valence images.
+This indicates that negative emotions are more easily evoked than positive ones and can be achieved through global edits to the source image.
+In addition to its creativity in generating positive-valence images, our EmoEditor also exhibits strong understanding and editing capabilities in generating negative emotions.
+In example(a), EmoEditor not only darkens the image’s tone to create a fearful atmosphere but also transforms the field into a desolate scene with weeds and branches, effectively evoking the emotion of fear.
+In example(c), it submerges the person, previously relaxing on a swim ring, beneath the pool water and alters his appearance to look frail and distressed, resembling a drowning victim, to evoke sadness.
+This highlights that EmoEditor can comprehend the content of the source image, align with its original emotional state, and combine it with the target emotion to perform specific edits, addressing both global and local aspects.
+S3.3 CrossTest Evaluation Results
+Our method employs an emotion predictor to guide emotional expression during image editing.
+However, using the same type of predictor for both generation and evaluation can introduce bias, as models may exploit specific predictor behaviors to inflate performance.
+To mitigate this, we split EmoSet into two disjoint subsets and trained two independent emotion predictors: one for generation and one for evaluation.
+As shown in Table S2, NST achieves the highest ESMI score.
+However, as illustrated in our visualization results, its outputs often feature emotion-colored overlays that obscure the original image content.
+These artifacts tend to inflate $S_{\text{emo}}$ because the emotion predictor is overly sensitive to low-level color cues rather than genuine emotional semantics.
+This exposes a key limitation of current evaluation practices and underscores the need for more semantically grounded emotion understanding.
+In contrast, results from our human study (Figure S9) show a clear preference for our method.
+Participants favored images where emotion-evoking regions were meaningfully altered, rather than simply covered with ambiguous color patches.
+This suggests that semantic coherence in emotional edits is more impactful and preferable than superficial visual modifications.
+Our method ranks second, only 2% below NST, while outperforming all other baselines.
+Unlike NST, it preserves visual semantics while achieving strong emotional alignment, demonstrating robust generalization and a more balanced trade-off between content fidelity and emotional expressiveness.
+ESMI(%) ↑ CT NST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
+CrossTest 28.39 45.35 32.61 39.38 38.50 26.54 28.67 38.50 34.09 43.95
+Table S2: Quantitative Comparison of Competing Methods under CrossTest.
+CrossTest refers to the evaluation scheme where images are generated using an emotion predictor trained on the EmoSet subset 1 and tested by an emotion predictor trained on the EmoSet subset 2.
+Bold indicates the best performance, while the second-best is underlined.
+Larger (↑) is better.
 S3.4 Limitations
-Figure S12 shows some failure cases.
-Our EmoEditor encounters difficulties in handling fine details of individuals, especially when their faces in the images are small.
-For example, in Row 1, although our EmoEditor attempts to make the boy appear sad, the boy’s face seems somewhat twisted and distorted.
-Moreover, significant disparities between source images and target emotions present obstacles, making it difficult to generate results aligning with the target emotion while preserving the structure of the source image.
-For example, in Row 2, the source image features a tiger, evoking the emotion of anger.
-In our EmoEditor’s attempt to transform it into awe, although the image’s structure is retained, it is altered to depict steep coastal rocks, changing the semantic content of the image.
-Figure S12: Failure Cases. The target emotion is highlighted in red and the source image is framed with green.
+While EmoEditor excels in quantitative performance and visual results, it encounters difficulties in handling fine details of individuals, especially when their faces in the images are small.
+For example, in Figure S13, although our EmoEditor attempts to make the boy appear sad, the boy’s face seems somewhat twisted and distorted.
+Moreover, although we use the 8-category Mikels model, real emotions are far more diverse and complex, and distinguishing positive emotions is often more challenging than negative ones.
+This is an area for future exploration.
+Figure S13: Failure Case. The target emotion is highlighted in red and the source image is framed in green.
+Acknowledgements
+This research is supported by the National Research Foundation, Singapore under its NRFF award NRF-NRFF15-2023-0001 and Mengmi Zhang’s Startup Grant from Nanyang Technological University.